### v1

# An Efficient and Adaptive Next Edit Suggestion Framework with Zero Human Instructions in IDEs

Abstract.
Code editing, including modifying, refactoring, and maintaining existing code, is the most frequent task in software development and has garnered significant attention from AI-powered tools. However, existing solutions that translate explicit natural language instructions into code edits face critical limitations, such as heavy reliance on human instruction input and high latency, which hinder their effective integration into a developer’s workflow.
We observe that developers’ habitual behaviors and coding objectives are often reflected in their historical editing patterns, making this data key to addressing existing limitations. To leverage these insights, we propose NES (Next Edit Suggestion), an LLM-driven code editing framework that delivers an instruction-free and low-latency experience. Built on a dual-model architecture and trained with our high-quality SFT and DAPO datasets, NES enhances productivity by understanding developer intent while optimizing inference to minimize latency.
NES is a scalable, industry-ready solution with a continuous Tab key interaction workflow, seamlessly adopted by a FinTech company with over 20,000 developers. Evaluations on real-world datasets show NES achieves 75.6% and 81.6% accuracy in two tasks of predicting next edit locations, alongside 91.36% ES and 27.7% EMR for intent-aligned edits, outperforming SOTA models. Our open-sourced SFT and DAPO datasets have been demonstrated to enhance the performance of open-source CodeLLMs. The demonstration of NES is available at https://youtu.be/yGoyYOe6fbY.
Code Edit, Reinforcement Learning, Large Language Models, Software Engineering
ccs: Software and its engineering
1. Introduction
With Large Language Models (LLMs) demonstrating remarkable capabilities in various aspects of software development (Di et al., 2023; Yang et al., 2025; Wang et al., 2024b; Guo et al., 2024; Liang et al., 2024; Rozière et al., 2023; Ding et al., 2024; Codefuse et al., 2025), AI-powered development tools such as GitHub Copilot (Copilot, 2023), Cursor (Cursor, 2024) and Zeta (Zed, 2025) are revolutionizing modern software development workflows by leveraging LLMs for code editing tasks (Wang et al., 2021; Zhang et al., 2022b; Gupta et al., 2023; Wei et al., 2024; Liu et al., 2024a). Empirical studies further highlight that code editing, which includes activities such as modifying (Li et al., 2024; Guo et al., 2025), refactoring, and maintaining existing code, is one of the most frequent and critical activities throughout the lifecycle of a software project (Nguyen et al., 2013). The importance of efficient and accurate code editing cannot be overstated, as it directly impacts the productivity and quality of software development.
Existing research efforts have made significant advancements, particularly in translating explicit natural language instructions into code edits. However, these approaches face critical limitations that hinder their effective integration into a developer’s workflow. On one hand, most existing methods rely heavily on human instruction input to maintain a coherent reference chain, disrupting programming flow and increasing cognitive load. Additionally, tools that fail to capture users’ habitual coding behavior often generate suggestions misaligned with their natural coding style. On the other hand, code editing tasks are inherently time-sensitive, with users expecting feedback within 1s (Card, 2018). Existing tools, however, rely on computationally intensive processes (e.g., search, natural language understanding), creating a tradeoff between low-latency and precise suggestions. This significant latency undermines real-time interaction and degrades user experience. Overcoming these limitations can undoubtedly enhance code editing in real-world development. However, it necessitates modifying the method of acquiring cues, which can pose significant challenges in accurately identifying the editing location and determining the content within the desired timeframe.
Our approach is based on an empirical observation: users’ habitual behaviors and coding objectives are often embedded within their historical editing patterns (e.g., repetitive refactoring actions or managing cross-file dependencies), which provide valuable contextual cues. Using historical editing patterns instead of relying on human instructions can minimize interruptions to the programming flow, maintain user-style consistency in suggestions, and enhance both the accuracy and efficiency of reference generation.
We propose NES (Next Edit Suggestion), an LLM-driven code editing framework designed for an instruction-free and low-latency (feel like zero-latency) experience. Comprising a synergistic dual-model architecture trained on the proposed Supervised Fine-Tuning (SFT) and Dynamic sAmpling Policy Optimization (DAPO) datasets, NES adapts to user habits, integrates seamlessly into workflows, and enhances productivity by implicitly learning developer intent.
  • NES-Location Model utilizes the developer’s historical editing patterns to predict the most probable next editing location, facilitating proactive and efficient navigation across the codebase.
  • NES-Edit Model analyzes the developer’s editing history, including refactoring patterns and API call sequences, to generate personalized and precise code modifications for the current editing region.
To improve user experience, NES leverages Prefix Caching (PC) (Zheng et al., 2023) and Speculative Decoding (SD) (Leviathan et al., 2022; Chen et al., 2023) techniques to achieve substantial inference speedups.
NES provides a scalable, industry-ready solution through its continuous Tab key interaction workflow, enabling seamless adoption in large-scale environments. Deployed in a FinTech company with 20,000+ developers, it handles thousands of daily code changes effortlessly. By allowing developers to perform complex tasks like refactoring or API call chains through simple "TabTabTab " actions, NES eliminates manual instructions, reduces cognitive load, and minimizes context-switching, delivering an efficient and immersive coding experience.
We evaluate NES using extensive real-world datasets from both open-source and industrial projects, highlighting its practical effectiveness. Experimental results show that the NES-Location Model achieves 75.7% accuracy in predicting the next edit location, while the NES-Edit Model delivers a 27.7% exact match rate (EMR) for generating intent-aligned code edits, outperforming several state-of-the-art (SOTA) models. Moreover, NES reduces inference time to under 450ms by optimizing model performance on Nvidia L20 GPUs. Finally, training on our proposed SFT and DAPO datasets results in notable improvements on Qwen2.5-Coder-7B and Seed-Coder-8B-Instrcut models.
Our key contributions are as follows.
  • We propose NES  [^1] NES mentioned in the paper, unless specifically noted otherwise, use the Qwen3-4B+SFT+DAPO model. See Section 5 for details., a dual-model, LLM-driven code editing framework that leverages historical editing patterns and operates in an instruction-free, low-latency paradigm, eliminating the need for explicit natural language human input.
  • We build the high-quality SFT and DAPO datasets [^2] We have submitted the datasets to GitHub and will make them publicly available soon at https://anonymous.link.for.review,which offer significant improvements in performance. These datasets have been instrumental in achieving accuracy and similarity scores that are dozens of times higher on code editing tasks across a variety of models.
  • We have successfully integrated and deployed NES on Nvidia L20 GPUs within a FinTech organization. Utilizing a comprehensive set of inference optimizations, NES has achieved low-latency performance, with response times under 450ms, thereby enhancing the user experience.
2. Motivating Example
This section motivates NES through a representative refactoring task in modern web development. For example, a software engineer enhancing UI components must add an aria-label property to all interactive buttons for improved screen reader accessibility (see Figure 2 for a common design pattern with a shared interface and multiple implementing components).
Figure 1. A motivating example of NES.
Figure 2. NES’s suggestion accepted by the developer with a single keystroke.
Without an adaptive assistance framework like NES, the developer must undertake a series of laborious steps:
  1. (1)
    Initial Modification: The developer updates the ButtonProps interface to include aria-label:string; property (Line 5).
  2. (2)
    Edit 1: The PrimaryButton is manually updated to de-structure the new prop and apply it to the <button> element (Line 8-9).
  3. (3)
    Edit 2: The same changes are repeated for the SecondaryButton component(Line 11-12).
  4. (4)
    Edit 3: The developer traces component usages, such as in the UserProfile component, and adds aria-label values to instances of <PrimaryButton/> (Line 22) and <SecondaryButton/> (Line 24).
The process involves manual navigation, repetitive edits, and significant cognitive effort to track changes. Although LLM-driven tools can assist with locating and making edits, instruction-based tools, such as CodeT5 (Wang et al., 2021), CoditT5 (Zhang et al., 2022b), Coeditor (Wei et al., 2024), and CoEdPilot (Liu et al., 2024a), disrupt the developer’s workflow by requiring precise commands and waiting for responses, introducing additional cognitive overhead. To address these challenges, an instruction-free tool is essential, which is where NES comes in.
NES streamlines this task by predicting the developer’s implicit intent based on their actions, enabling a seamless, fluid interaction.
  1. (1)
    Intent Seeding: The developer initiates the process by intentionally adding the aria-label: string; property to the ButtonProps interface, signaling a cue for the NES framework.
  2. (2)
    Proactive Edit Suggestion (NES-Edit): When the cursor moves near the PrimaryButton definition, the NES-Edit model detects the dependency change and familiar refactoring pattern, proactively suggesting updates (see Figure 2) to the component’s signature and implementation.
  3. (3)
    Predictive Navigation and Chained Editing (NES-Location + NES-Edit): The NES-Location model predicts the next logical change at the SecondaryButton component, guiding the cursor to its definition. Upon arrival, the NES-Edit model suggests the required refactoring, accepted with a single Tab press.
  4. (4)
    Cascading Assistance: The framework continues predicting and assisting, leading to call sites in UserProfile. Here, NES-Edit suggests context-specific values (e.g., aria-label="Save Changes") for the new property, with each step accepted via Tab.
A tedious, error-prone manual process is streamlined into an efficient "TabTabTab " workflow. Powered by high-quality datasets, NES predicts both the location (NES-Location) and content (NES-Edit) of edits without requiring explicit instructions. This reduces cognitive load, minimizes context switching, accelerates inference and delivers an immersive development experience.
3. Methodology
Figure 3. Illustration of the NES framework’s workflow.
3.1. An Overview of NES Framework
Figure 3 provides a detailed illustration of the NES framework’s workflow. The framework’s functionality is initiated by the developer’s initial editing action, subsequently entering a dynamic interactive cycle:
  • Next Location Suggestion: The NES-Location Model, leveraging real-time updated historical editing trajectories ①, predicts the developer’s potential next edit location. Subsequently, the framework presents this predictive suggestion to the developer, who then has the option to either rapidly navigate to the suggested position by pressing the Tab key or to disregard the suggestion ②.
  • Code Edit Suggestion: Thereafter, the NES-Edit Model, integrating historical editing patterns with the current editing window’s context, generates real-time code editing suggestions ③. Developers have the flexibility to either accept these suggestions by pressing the Tab key or continue with independent edits ④, while also updating the edit history ⑤.
  • Dynamic loop: Upon the developer’s adoption of an editing suggestion, the framework promptly initiates the next round of next location suggestion, establishing a continuous intelligent assistance loop. Ideally, developers can significantly enhance their coding efficiency and quality by consistently accepting NES framework’s suggestions by pressing the Tab key ⑥.
To implement this framework, we focused on four key areas:
  • Incremental Difference Detection: We developed a real-time code change monitor based on incremental computation. This framework accurately captures each developer editing operation, thereby constructing a comprehensive and detailed historical editing trajectory. These trajectory data serve as crucial inputs for subsequent model training and prediction processes.
  • High-Quality Training Dataset Construction: Utilizing the collected authentic historical editing trajectories, we constructed a high-quality training dataset. This dataset encompasses not only sequences of editing operations but also contextual information for each edit. Therefore, this dataset can enabling the training of both the NES-Location Model and the NES-Edit Model with nuanced understanding of editing patterns and their surrounding code environments.
  • Model Training: To endow NES with the capability for location and code editing suggestion, we designed a two-stage training methodology. Initially, both the NES-Location and NES-Edit models undergo SFT (Ouyang et al., 2022) on our large-scale historical editing datasets, enabling them to learn the fundamental patterns of code modification. Subsequently, we employ Decoupled Clip and DAPO (Yu et al., 2025), a reinforcement learning technique using high-quality preference data, to further refine the models. This second stage aligns their behavior more closely with real-world developer intent and utility, significantly boosting the practical value of NES.
  • Model Inference Optimization: Acknowledging the constraints of inference resources in practical application environments and the stringent low-latency requirements inherent in real-time editing suggestion, we conducted in-depth optimization of the model inference process, and achieved real-time responsiveness to developer editing behaviors, ensuring the fluidity and practicality of the NES framework.
We present the details of these critical components in the following sections, respectively.
3.2. Incremental Difference Detection
We developed an efficient incremental difference detection mechanism built upon our AI-powered IDE, designed to accurately capture real-world code editing trajectories of developers during real-world programming activities. As detailed in Algorithm  1, it operates via two steps: incremental difference calculation and difference merging, where "Overlap" denotes intersecting regions of two editing operations.
3.2.1. Incremental Difference Calculation
To ensure real-time responsiveness, we replaced conventional file-wide difference calculation methods with an incremental approach. The core principle is to narrow the computational scope from the entire file to only the code segments actively being modified by the developer. By focusing solely on code segments actively modified by the user, our AI-powered IDE captures edit regions in real-time and applies a sequence-matching method to these localized segments, minimizing computational overhead and enabling instant change detection.
To provide the model with precise and unambiguous context, we designed the custom NES diff format. As illustrated in Table 1, our format enriches the standard diff by prefixing every line—whether added (+), deleted (-), or unchanged—with its absolute line number. The NES diff format provides clearer code modification and line number change information for the model, directly improving the accuracy of location and code editing tasks.
3.2.2. Instant Difference Merge
The high-frequency differential computation in our framework results in the acquisition of extensive granular editing histories. Utilizing such fragmented data for model training would potentially lead to the generation of incomplete and atomized suggestions, necessitating frequent user interactions to adopt multiple fragmentary recommendations. To address this, we implemented instant merging of overlapping differences, using developer-edited code blocks as units to create cohesive recommendations aligned with real-world editing behaviors.
difflib format NES diff format
@@ -1,3 +1,3 @@ -def Hello() +def GoodBye() print("Say") - print("Hello") + print("GoodBye") 1-| def Hello() 1+| def GoodBye() 2 | print("Say") 3-| print("Hello") 3+| print("GoodBye")
Table 1. An example comparison between the diff formats of difflib and NES highlights that NES’s format offers higher information density with precise line numbers.
Algorithm 1 Incremental Difference Detecting
A stream of edit events $E_{stream}$, where each event $e=(C_{pre},C_{post})\in E_{stream}$ encompasses the pre- and post-modification content of code snippets.
A list of merged historical edits =$\mathcal{H}=\{\Delta_{1},\Delta_{2},\dots\}$.
Initialize $\mathcal{H}\leftarrow\emptyset$ and $\Delta_{active}\leftarrow\text{null}$.
for each incoming edit event $(C_{pre},C_{post})$ in $E_{stream}$ do
  $\Delta_{new}\leftarrow\text{Diff}(C_{pre},C_{post})$
  if $\Delta_{active}$ is null then
    $\Delta_{active}\leftarrow\Delta_{new}$
  else if $\text{Overlap}(\Delta_{active},\Delta_{new})$ then
    $\Delta_{active}\leftarrow\text{Merge}(\Delta_{active},\Delta_{new})$
  else
    Add $\Delta_{active}$ to $\mathcal{H}$
    $\Delta_{active}\leftarrow\Delta_{new}$
  end if
end for
After the stream ends, add the last non-null $\Delta_{active}$ to $\mathcal{H}$.
return the list of historical trajectories $\mathcal{H}$.
Figure 4. Training dataset collection process
3.3. Training Dataset Construction
Building a high-quality training dataset is key to NES’s success. Using developer editing trajectories (Section 3.2), we convert raw data into structured instances for our dual-model architecture through two stages: instance formulation and relevance-based filtering.
3.3.1. Formulation of Sequential Edit Instances.
Our process begins by segmenting the continuous stream of code modifications into discrete, sequential training instances. As illustrated in Figure 4, we consider any two adjacent timestamps, $T$ and $T+1$, where a developer’s action (e.g., an edit) creates a state transition. Each transition yields a data tuple containing:
  • The pre-edit code state ($C_{T}$).
  • The historical editing trajectory up to that point ($H_{T}$).
  • The ground-truth edit, comprising the location ($L_{\text{gt}}$) and content ($E_{\text{gt}}$) of the modification performed between $T$ and $T+1$.
From this tuple, we formulate training samples:
  • For the NES-Location Model: The input is the pair ($C_{T}$, $H_{T}$), and the target label is the subsequent edit location, $L_{\text{gt}}$.
  • For the NES-Edit Model: The input consists of ($C_{T}$, $H_{T}$) and the ground-truth edit location $L_{\text{gt}}$. The target label is the new code content, $E_{\text{gt}}$.
In this manner, we established our preliminary training dataset.
3.3.2. Relevance-Based Filtering and Negative Sampling.
A significant challenge with raw trajectory data is the prevalence of spurious correlations. Developers’ edit sequences are not always causally linked; a developer might complete one task and immediately start an unrelated one. Training on such uncorrelated sequences could lead the model to learn erroneous associations, resulting in disruptive and irrelevant suggestions.
To mitigate this, we introduce a crucial data refinement stage. We employ an LLM as a relevance filter to analyze the semantic and causal relationship between the historical edits ($H_{T}$) and the current edit ($L_{\text{gt}}$, $E_{\text{gt}}$). Based on the LLM’s assessment, each instance is processed as follows:
  • Relevant Edits: If the LLM filter determines that the current edit is a logical and predictable continuation of the historical trajectory, the instance is treated as a modification (-do) sample. The original ground-truth edit is retained, teaching the model that the history provides a valid basis to predict the corresponding modification or navigation.
  • Irrelevant Edits: Conversely, if the filter assesses the edit as uncorrelated with the history, it implies that the developer’s next action cannot be reliably predicted from the preceding edits. In this case, the instance is repurposed as a preservation (-keep) sample. This crucial step teaches the model to refrain from making a suggestion when the developer’s intent is not inferable from the available history.
This filtering and relabeling process transforms the preliminary training datacset into a final, optimized training set. This refined dataset not only enhances the quality of modification-oriented samples but, critically, also incorporates explicit negative samples. This dual focus is essential for training a model that is both proactive in its assistance and discerning enough to avoid generating noise.
Finally, it is essential to elucidate the language distribution of our dataset. Due to the characteristics of our data source developers, the training data exclusively consists of two frontend languages: TypeScript and TypeScriptReact. Nevertheless, as demonstrated in Section 4, our model demonstrates a certain degree of generalization capability on backend languages such as Java and Python.
3.4. Model Training
Our training strategy for the NES-Location and NES-Edit models follows a two-stage process to optimize performance and align with developer preferences. We adapt three pre-trained, small-sized code LLMs for our tasks: Qwen3-4B, Qwen2.5-Coder-7B, and Seed-Coder-8B-Instruct. The training process involves: (1) SFT to establish foundational task capabilities, and (2) Reinforcement Learning with DAPO to further refine the model’s behavior based on task-specific rewards.
3.4.1. Stage 1: SFT
The initial stage uses SFT to adapt the base code LLMs to our specific tasks. The goal of this stage is to teach the models the fundamental input-output formats required for next edit suggestion (as shown in Figure 5). We use our large-scale historical editing datasets (~200k samples for NES-Edit, ~60k for NES-Location) for this purpose. During SFT, the model learns to process a structured input, which includes the $C_{T}$, $H_{T}$, and the $L_{\text{gt}}$ (for the NES-Edit model), to generate a plausible output—either $L_{T+1}$ or $E_{T+1}$. This step establishes a robust baseline model with a solid understanding of code context and common editing patterns.
Figure 5. Input-Output Format
3.4.2. Stage 2: Reinforcement Learning with DAPO
SFT teaches patterns, but lacks task-specific optimization. RL with DAPO refines the model by rewarding correct and accurate outputs. In the RL phase, a custom reward function evaluates generated answers ($L_{\text{gen}}$ or $E_{\text{gen}}$) against ground truth based on format correctness and content accuracy, guiding the model toward well-formed and precise outputs.
Reward Function for NES-Location Model
In contrast, the task of the NES-Location model is to predict a discrete, precise location. In this context, the concept of partial credit is less applicable. Therefore, the reward function $R_{\text{Location}}$ is a straightforward binary signal based purely on accuracy:
(1) $R_{\text{Location}}(L_{\text{gen}},L_{\text{gt}})=\begin{cases}1.0&\text{if }L_{\text{gen}}\text{ = }L_{\text{gt}}\\ -1.0&\text{otherwise}\end{cases}$
Reward Function for NES-Edit Model. The reward function for the NES-Edit model is designed to prioritize perfect accuracy while also assigning partial credit to semantically similar, potentially useful suggestions. This is achieved through a hierarchical reward structure based on Exact Match Rate (EMR) and Edit Similarity (ES). The reward $R_{\text{Edit}}$ is defined as:
(2) $R_{\text{Edit}}(E_{\text{gen}},E_{\text{gt}})=\begin{cases}1.0&\text{if }E_{\text{gen}}\text{ = }E_{\text{gt}}\\ 0.5\times\text{ES}(E_{\text{gen}},E_{\text{gt}})&\text{if }\text{ES}(E_{\text{gen}},E_{\text{gt}})>0.5\\ -1.0&\text{otherwise}\end{cases}$
A reward of 1.0 is granted for an exact match, strongly encouraging the model to pursue perfect accuracy. If the generated edit is not an exact match but is substantially similar to the ground truth (ES > 0.5), a proportional reward is assigned. The scaling factor of 0.5 ensures this partial credit is strictly less than the reward for a perfect solution, while still guiding the model to generate semantically relevant code. A penalty of -1.0 is given for suggestions that are neither correct nor sufficiently similar, preventing the model from learning to produce irrelevant outputs.
This clear and unambiguous reward signal effectively trains the model to maximize its accuracy in predicting the developer’s next point of focus. By applying DAPO with these carefully designed reward functions, we enhance the models’ practical utility beyond what can be achieved with SFT alone.
3.5. Optimization of Model Inference
The NES framework demands low latency to ensure a seamless developer experience. To tackle this, we implemented various inference optimization techniques during deployment:
  • Efficient Model Deployment Based on vLLM: Used the vLLM framework, leveraging Paged Attention and Dynamic Batching to optimize LLM inference for performance, efficiency, and usability.
  • Speculative Decoding: Applied Speculative Decoding (Leviathan et al., 2022) to pre-generate candidate code segments with a lightweight n-gram model, enabling the NES-Edit Model to verify or refine them, reducing computational overhead.
  • Prefix Caching: Reused intermediate results by caching Key-Value (KV) Cache from historical requests, improving inference efficiency for repetitive prefixes.
4. Evaluation
We evaluate NES by addressing the following research questions:
  • RQ1 Next Location Prediction: To what extent can NES accurately predict the position of the next edit?
  • RQ2 Code Edit Generation: How effectively can NES generate precise edit suggestions?
  • RQ3 Model Performance: Does NES demonstrate sufficient computational efficiency to provide low-latency suggestions in interactive development environments?
4.1. Experiment Setup
4.1.1. Benchmarks
To evaluate the performance of NES, we construct a benchmark of 4 programming languages (i.e., Java, Python, TypeScript and TypeScriptReact) from 1000 open-source and industrial projects. For the open-source projects, we selected the top 100 repositories on GitHub based on star count, while the industrial projects were chosen based on their update frequency. Following the same methodology used for training data, we constructed a unified test dataset comprising 1,000 samples per language. This dataset is systematically divided into two distinct scenarios: Modification and Preservation.
  • Modification (-do): 500 samples per language where subsequent edits occur in a different location, testing the model’s predictive navigation capability.
  • Preservation (-keep): 500 samples per language where edits remain in the same location or terminate, assessing the model’s ability to avoid unnecessary navigation suggestions.
4.1.2. Metrics
For RQ1, which evaluates the prediction of the next editing position, the primary metric used is Accuracy (ACC), measuring whether the predicted edit location matches the ground truth.
For RQ2, which evaluates code-editing performance, the three key metrics are: Edit Similarity (ES) for semantic alignment between the generated and target edits; Exact Match Rate (EMR) for syntactic precision of the generated edits for the modification task; and Accuracy (ACC) for correctly predict preservation.
For RQ3, which investigates model inference speed, the performance metrics are input and output tokens/s speed.
4.1.3. Baseline Models
We compare NES to the following models.
  • Qwen (Yang et al., 2025), developed by Alibaba Cloud, is a renowned LLM, while Qwen-Coder (Yang et al., 2024) specializes in programming and software development tasks.
  • Claude (Anthropic, [n. d.]), created by Anthropic, has strong capabilities in coding and can assist with a wide range of programming tasks, including writing, debugging, and explaining code.
  • Seed-Coder (Seed et al., 2025), recently released by ByteDance, is a lightweight yet powerful open-source code-focused large language model (LLM) that delivers strong performance across a variety of coding benchmarks.
  • Zeta (Zed, 2025), recently introduced by Zed, is an advanced tool designed to provide instruction-free edit generation capabilities. Zeta intuitively generates context-aware code or text modifications based solely on the input provided.
We categorize these models into two groups based on their size. The models with relatively smaller sizes include Qwen3-4B, Qwen2.5-Coder-7B, and Seed-Coder-8B-Instruct. These models will undergo supervised fine-tuning (SFT) using the NES SFT and DAPO datasets. Zeta-7B, however, was not subjected to additional training, as it is already built on Qwen2.5-Coder-7B and incorporates Zeta’s own training data. Instead, we use Zeta-7B as a baseline.
The models with relatively larger sizes including Qwen3-32B, Qwen2.5-Coder-32B-Instruct, Claude3, and Claude4. These were not trained further due to the following reasons: a. Large models are challenging to meet the low-latency requirements in code editing scenarios. b. Training large models is resource-intensive and consumes significant computational resources.
We did not select instruction-based tools such as CodeT5 (Wang et al., 2021), CoditT5 (Zhang et al., 2022b), Coeditor (Wei et al., 2024), and CoEdPilot (Liu et al., 2024a) as baselines, as these tools rely on explicit human instructions to operate. Since our evaluation environment and benchmarks do not account for such input, including these tools would lead to an unfair comparison. However, based on feedback from real-world developers, our approach provides an editing tool that is at least as effective, if not superior, in delivering practical assistance to users.
4.1.4. Settings
In our testing environment, we utilized a single NVIDIA A100-80GB GPU for evaluation purposes. For the production deployment in our online real-world environment, we implemented a configuration of NVIDIA L20-48GB GPUs. Both setups were based on vLLM version 0.8.5.
4.2. Results
Model TypeScript TypeScriptReact Java Python Average
do keep do keep do keep do keep do keep
Qwen2.5-Coder-7B 8.7% 2.2% 9.8% 6.8% 10.3% 1.9% 11.8% 3.7% 10.1% 3.7%
Seed-Coder-8B-Instruct 8.0% 1.9% 9.0% 3.7% 6.5% 1.8% 13.5% 4.1% 9.3% 2.9%
Qwen3-4B 5.6% 1.8% 4.4% 2.0% 7.7% 0.8% 8.7% 3.0% 6.6% 1.9%
Qwen3-32B 51.6% 6.7% 46.4% 3.1% 64.2% 1.0% 63.0% 2.9% 56.3% 3.4%
Qwen2.5-Coder-32B-Instruct 15.6% 15.5% 12.8% 17.1% 12.9% 15.7% 27.2% 18.2% 17.1% 16.6%
Claude3 57.2% 5.4% 54.2% 3.8% 63.6% 21.0% 66.4% 3.4% 60.4% 8.4%
Claude4 65.2% 8.0% 65.0% 10.0% 60.0% 15.6% 77.6% 4.7% 67.0% 9.6%
Qwen2.5-Coder-7B+SFT 62.4% 94.6% 69.3% 97.0% 51.3% 95.3% 68.4% 89.3% 62.8% 94.0%
Seed-Coder-8B-Instruct+SFT 73.2% 94.6% 75.7% 96.6% 64.1% 90.4% 77.4% 94.8% 72.6% 94.1%
Qwen3-4B+SFT 74.6% 90.5% 76.9% 91.6% 64.1% 89.3% 74.6% 86.0% 72.6% 89.4%
Qwen2.5-Coder-7B+SFT+DAPO 66.5% 85.3% 71.7% 89.4% 53.6% 86.3% 70.4% 85.1% 65.6% 86.5%
Seed-Coder-8B-Instruct+SFT+DAPO 77.6% 85.1% 78.9% 87.8% 73.9% 83.6% 80.4% 83.6% 77.7% 85.0%
Qwen3-4B+SFT+DAPO 76.5% 82.3% 77.3% 86.2% 71.8% 80.1% 76.9% 77.8% 75.6% 81.6%
Table 2. Accuracy comparison for location tasks (-do and -keep) across models. “model+SFT” and “model+SFT+DAPO” indicate training on the base model with the NES SFT dataset alone and on both the SFT and DAPO datasets.
Max Edit History TypeScript TypeScriptReact Java Python Average
do keep do keep do keep do keep do keep
1 73.7% 84.1% 73.9% 89.4% 69.2% 84.7% 75.9% 84.9% 73.2% 85.8%
3 74.6% 90.5% 76.9% 91.6% 64.1% 89.3% 74.6% 86.0% 72.6% 89.4%
5 68.7% 80.4% 70.5% 82.6% 56.4% 78.8% 73.2% 76.1% 67.2% 79.5%
7 68.3% 79.8% 70.6% 81.8% 56.4% 78.8% 74.1% 76.1% 67.4% 79.1%
9 67.6% 82.2% 71.4% 86.2% 60.0% 80.1% 72.1% 75.7% 67.8% 81.0%
Table 3. Comparison of location prediction performance with varying maximum edit history lengths on the Qwen3-4B+SFT model
4.2.1. RQ1 NES-Location:
To what extent can NES accurately predict the location of the next edit?
Table 2 demonstrates the accuracy of the NES-Location model in two distinct scenarios: modification (-do) and preservation (-keep). The terms “model+SFT” and “model+SFT+DAPO” indicate training on the base model with the NES SFT dataset alone and on both the NES SFT and DAPO datasets, respectively.
Impact of SFT: Before post-training, the best base model, Claude 4, achieves 67.0% and 9.6% average accuracy on the -do and -keep tasks, respectively. Models with small size, such as Qwen2.5-Coder-7B, perform poorly, with only 10.1% and 3.7% on the -do and -keep tasks, respectively. NES SFT dataset significantly improves models with small size. For -do, Qwen2.5-Coder-7B jumps from 10.1% to 62.8%, Seed-Coder-8B-Instruc from 9.3% to 72.6%, and Qwen3-4B from 6.6% to 72.6%. For -keep, Qwen2.5-Coder-7B rises from 3.7% to 94.0%, Seed-Coder-8B-Instruc from 2.9% to 94.1%, and Qwen3-4B from 1.9% to 89.4%.
Due to the limits of low-latency requirements and training resource constraints, large models were not trained on the SFT dataset. Additionally, Zeta does not provide navigation functionality, so its results are excluded from Table 2.
RQ1 Answer about SFT: The NES SFT dataset significantly boosts model performance in predicting the next editing position, improving Qwen3-4B’s accuracy from 6.6% to 72.6% on the -do task and from 1.9% to 89.4% on the -keep task.
Impact of DAPO: Models with smaller sizes trained on the DAPO dataset achieve slight improvements in average navigation prediction (-do) accuracy: Qwen2.5-Coder-7B improves from 62.8% to 65.6%, Seed-Coder-8B-Instruc increases from 72.6% to 77.7%, and Qwen3-4B rises from 72.6% to 75.6%. However, on the -keep task, Qwen2.5-Coder-7B declines from 94.0% to 86.5%, Seed-Coder-8B-Instruct declines from 94.1% to 85.0%, and Qwen3-4B drops from 89.4% to 81.6%.
The DAPO-trained model tends to suggest location changes more frequently. Because changing is more frequently in the real-world development practices, We accept the minor drop in -keep performance in favor of improved -do accuracy. The NES-Location Model aims to streamline navigation, so we’ve adopted DAPO, prioritizing the higher success rate in predicting necessary jumps over avoiding suggestions, as it better supports developers’ workflow.
RQ1 Answer about DAPO: The DAPO-trained model tends to suggest location changes more frequently, leading to a slight drop in -keep task accuracy. However, we have decided to adopt the DAPO-trained NES-Location model, which demonstrates better performance on the -do task, as location changes are more prevalent in real-world development practices.
Impact of Edit History Length: Table 3 demonstrates the impact of historical context length (Max Edit History) on Qwen3-4B+SFT. A history length of 3 achieves the best balance, with 72.6% -do average accuracy and 89.4% -keep average accuracy. Short histories (e.g., 1) lack sufficient context for complex intents, while longer histories (e.g., 5+) introduce irrelevant noise, hindering accuracy.
RQ1 Answer about Edit History Length: Short histories may lack context, while long histories can introduce noise, both hindering accuracy. The results show that a history length of 3 is optimal.
Model TypeScript TypeScriptReact Java Python Average
ES / EMR ES / EMR ES / EMR ES / EMR ES / EMR
Qwen2.5-Coder-7B 72.17/12.4% 69.52/6.4% 64.22/4.6% 72.52/8.8% 69.61/8.1%
Seed-Coder-8B-Instrcut 71.30/14.9% 69.35/7.0% 69.91/8.8% 70.06/9.8% 70.16/10.1%
Qwen3-4B 76.77/6.4% 80.46/13.5% 73.32/6.6% 77.23/7.0% 76.94/8.5%
Qwen3-32B 79.32/16.7% 76.81/11.4% 75.66/13.8% 79.01/15.2% 77.70/14.3%
Qwen2.5-Coder-32B-Instruct 81.33/19.6% 81.34/13.2% 76.39/14.0% 79.80/19.0% 79.72/16.5%
Zeta 86.22/17.3% 86.29/9.2% 86.34/15.0% 84.90/8.8% 85.94/12.6%
Claude3 84.81/24.3% 89.0/19.2% 82.81/18.8% 81.72/22.8% 84.59/21.3%
Claude4 85.37/25.7% 85.56/19.6% 83.40/22.4% 83.38/24.8% 84.43/23.1%
Qwen2.5-Coder-7B+SFT 92.49/28.8% 92.63/28.8% 89.17/19.6% 89.61/22.6% 90.98/25.0%
Seed-Coder-8B-Instrcut+SFT 92.29/30.2% 92.85/34.0% 89.55/20.4% 89.54/22.0% 91.06/26.7%
Qwen3-4B+SFT 92.03/27.0% 90.69/29.6% 89.29/20.2% 89.30/21.6% 90.33/24.6%
Qwen2.5-Coder-7B+SFT+DAPO 91.82/32.4% 92.67/34.4% 89.99/21.0% 90.03/25.4% 91.13/28.3%
Seed-Coder-8B-Instrcut+SFT+DAPO 91.90/31.4% 93.21/34.2% 90.15/23.6% 90.08/26.0% 91.34/28.8%
Qwen3-4B+SFT+DAPO 92.77/30.6% 93.28/34.2% 89.97/21.4% 89.41/24.6% 91.36/27.7%
Table 4. Similarity comparison for modification tasks (-do). “model+SFT” and “model+SFT+DAPO” indicate training on the base model with the NES SFT dataset alone and on both the SFT and DAPO datasets.
Model TypeScript TypeScriptReact Java Python Average
Qwen2.5-Coder-7B 11.6% 8.0% 4.6% 8.8% 8.3%
Seed-Coder-8B-Instrcut 2.6% 3.2% 2.0% 3.8% 2.9%
Qwen3-4B 35.2% 37.4% 28.8% 44.8% 36.6%
Qwen3-32B 14.0% 21.2% 9.4% 18.8% 15.9%
Qwen2.5-Coder-32B-Instruct 11.0% 11.0% 7.2% 24.2% 13.4%
Zeta 40.0% 35.4% 43.2% 44.2% 40.7%
Claude3 21.0% 17.6% 24.0% 19.4% 20.5%
Claude4 15.6% 15.2% 11.0% 17.0% 14.7%
Qwen2.5-Coder-7B+SFT 94.4% 90.2% 87.4% 92.0% 91.0%
Seed-Coder-8B-Instrcut+SFT 92.2% 86.0% 88.6% 89.8% 89.2%
Qwen3-4B+SFT 92.4% 84.8% 88.0% 88.8% 88.5%
Qwen2.5-Coder-7B+SFT+DAPO 84.8% 85.2% 80.8% 86.2% 84.3%
Seed-Coder-8B-Instrcut+SFT+DAPO 85.2% 80.0% 78.2% 83.4% 81.7%
Qwen3-4B+SFT+DAPO 86.2% 88.8% 82.0% 88.2% 86.3%
Table 5. Accuracy comparison for preservation tasks (-keep). “model+SFT” and “model+SFT+DAPO” indicate training on the base model with the NES SFT dataset alone and on both the SFT and DAPO datasets.
Ratio of TypeScript TypeScriptReact Java Python Average
Unmodified do keep do keep do keep do keep do keep
Data ES / EMR Acc ES / EMR Acc ES / EMR Acc ES / EMR Acc ES / EMR Acc
10% 91.44/26.7% 92.2% 91.22/21.2% 87.0% 89.03/18.8% 87.6% 88.95/21.6% 89.6% 90.16/22.1% 89.1%
20% 91.83/26.1% 93.4% 91.36/20.6% 89.6% 89.06/16.4% 90.2% 89.05/20.4% 92.2% 90.33/20.9% 91.4%
30% 91.73/24.9% 94.8% 90.81/20.4% 90.8% 88.11/15.4% 89.8% 89.19/19.6% 92.6% 89.96/20.1% 92.0%
40% 91.55/24.1% 96.4% 91.20/20.4% 92.0% 89.17/17.2% 91.6% 89.22/19.6% 94.0% 90.28/20.33% 93.5%
50% 91.61/24.3% 96.4% 90.87/18.6% 94.6% 88.72/14.2% 93.9% 89.40/19.2% 95.4% 90.15/19.1% 95.1%
60% 91.89/21.4% 98.6% 91.01/20.0% 96.0% 88.73/12.2% 95.4% 88.92/15.6% 95.8% 90.14/17.3% 96.5%
Table 6. Similarity comparison with different ratios of unmodified data
Max TypeScript TypeScriptReact Java Python Average
Edit do keep do keep do keep do keep do keep
History ES / EMR Acc ES / EMR Acc ES / EMR Acc ES / EMR Acc ES / EMR Acc
1 96.81/21.7% 93.8% 91.35/21% 85.4% 88.88/17.0% 88.4% 89.16/19.8% 91.2% 91.55/19.9% 89.7%
3 92.03/27.0% 92.4% 90.69/29.6% 84.8% 89.29/20.2% 88.0% 89.30/21.6% 88.8% 90.33/24.6% 88.5%
5 91.83/26.1% 93.4% 91.36/20.6% 89.6% 89.06/16.4% 90.2% 89.05/20.4% 92.2% 90.33/20.9% 91.4%
7 92.49/29.4% 95.0% 91.43/23.0% 90.8% 88.13/17.6% 89.2% 88.06/21.2% 91.6% 90.03/22.8% 91.7%
9 91.61/29.2% 93.0% 91.07/23.0% 87.8% 89.06/16.4% 90.2% 89.05/20.4% 92.2% 90.20/22.3% 90.8%
Table 7. Comparison of code edit generation with varying maximum edit history lengths on the Qwen3-4B+SFT model
4.2.2. RQ2 Code Edit Generation:
How effectively can NES generate precise edit suggestions?
Impact of SFT: Table 4 and Table 5 present the performance of the NES-Edit Model across four programming languages: Java, Python, TypeScript, and TypeScript React. In the -do task, Zeta achieves the highest ES score of 85.94%, while Claude4 records the best EMR score of 23.1%. Following SFT fine-tuning, smaller models demonstrate improved ES and EMR scores. For instance, Qwen2.5-Coder-7B+SFT achieves an ES of 90.98% and EMR of 25.0%, Seed-Coder-8B-Instruct+SFT achieves an ES of 91.06% and EMR of 26.7%, and Qwen3-4B+SFT achieves an ES of 90.33% and EMR of 24.6%. In the -keep task, Zeta achieves the highest initial accuracy of 40.7%. After SFT fine-tuning, smaller models show significant improvements in accuracy, with Qwen2.5-Coder-7B+SFT reaching 91.0%, Seed-Coder-8B-Instruct+SFT achieving 89.2%, and Qwen3-4B+SFT attaining 88.5%.
RQ2 Answer about SFT: The NES SFT dataset greatly improves model performance in code edit generation.
Impact of DAPO: Smaller models trained on the DAPO dataset achieve moderate improvements in average code edit generation (-do) similarity scores. For example, Qwen2.5-Coder-7B improves from an ES of 90.98% to 91.13% and an EMR of 25.0% to 28.3%; Seed-Coder-8B-Instruct increases from an ES of 91.06% to 91.34% and an EMR of 26.7% to 28.8%; and Qwen3-4B improves from an ES of 90.33% to 91.36% and an EMR of 24.6% to 27.7%.
For the -keep task, Qwen2.5-Coder-7B experiences a decline in accuracy from 91.0% to 84.3%, Seed-Coder-8B-Instruct sees an decrease from 89.2% to 81.7%, and Qwen3-4B drops from 88.5% to 86.3%.
The slight decrease in performance on the -keep task can be attributed to the same reason as with NES-Location: the DAPO-trained model tends to suggest location changes more frequently. Nevertheless, we deploy the DAPO-trained model in real-world scenarios due to its alignment with high-frequency practical use cases.
RQ2 Answer about DAPO: The DAPO-trained model demonstrates improved similarity scores for the -do task, yet shows a minor drop in accuracy for the -keep task. However, we have chosen to adopt the model trained with DAPO, as it better aligns with the high-frequency practices observed in real-world software development.
Impact of Unmodified Data: We investigated the impact of varying proportions of unmodified data in the training set, analyzing ratios from 10% to 60% (Table 6). The results show that higher proportions of unmodified data improve performance on the -keep test sets, while leading to a decline in performance on the -do test sets. After thorough analysis, a 20% proportion of unmodified data was determined to be optimal. As a result, we selected the 20% unmodified data ratio as the standard configuration for all subsequent experiments.
RQ2 Answer about Unmodified Data: As a result, we selected the 20% unmodified data ratio as the standard configuration for all subsequent experiments.
Impact of Edit History Length: Table 7 illustrates the effect of varying the length of historical context (Max Edit History) on the performance of Qwen3-4B+SFT. Optimal results are achieved with a history length of 3, yielding a -do average similarity score of ES 90.33% and EMR 24.6%. Conversely, a history length of 7 results in the highest -keep average accuracy of 88.5%. This suggests that generating better code edits may require a more extensive historical trajectory.
RQ2 Answer about Edit History Length: Generating better code edits may require a more extensive historical trajectory.
4.2.3. RQ3 Model Performance:
Does NES demonstrate sufficient computational efficiency to provide low-latency suggestions in interactive development environments?
Table 8 presents the results of our inference speed analysis. The data reveals that both Prefix Caching (PC) and Speculative Decoding (SD) significantly enhance inference speed, with Speculative Decoding providing the largest improvement. This aligns with our initial hypothesis, confirming the suitability of Speculative Decoding for our specific use case.
When compared to Zeta, a quantized 7B model with the same inference optimizations, NES achieves superior input throughput, while maintaining equivalent output throughput. In our online environment, NES ultimately attained an average inference latency of 450ms, effectively reaching real-time processing capabilities and successfully meeting our performance objectives.
Model configurations input tokens/s output tokens/s
NES-Edit+vLLM 4800 71
NES-Edit+vLLM+PC 5750 85
NES-Edit+vLLM+PC+SD 8500 125
Zeta+vLLM+PC+SD 5400 125
Table 8. Comparison of input and output token processing speeds for different model configurations
RQ3 Answer: The Speculative Decoding (SD) optimization delivers the most significant improvement in inference input/output throughput, enabling NES to achieve an average inference latency of 450ms.
5. Design of Industrial Tools
In transitioning the NES framework from a research prototype to a scalable, industrial-grade solution, we faced the challenge of balancing model performance, inference latency, and deployment cost. To meet the stringent responsiveness demands of real-time code editing, we chose the Qwen3-4B model as the core engine for NES, post-train it with our SFT and DAPO datasets (see Section 3.3) for production deployment. This decision was based on several key considerations:
  • Trade-off between Model Size and Latency: Larger scale models, like 8B-parameter ones, have high computational needs and slower inference, making them unsuitable for real-time code editing’s low-latency requirements. The Qwen3-4B model, being smaller, offers faster inference, aligning better with our goal of achieving a "zero-latency" user experience in NES.
  • Excellent Cost-Performance Ratio after DAPO Training: As detailed in Section 4, our two-stage training process greatly improved Qwen3-4B’s performance. The Qwen3-4B+SFT+DAPO model ranks among the best small-scale models and rivals larger ones. While DAPO slightly reduces performance on the -keep task, this trade-off is acceptable due to its superior results on the more frequent practical -do task, validating our strategy of leveraging high-quality data and advanced training to balance performance and efficiency in a compact model.
  • Feasibility for Large-Scale Deployment: In industrial applications, ensuring service stability and scalability is critical. We deployed the Qwen3-4B+SFT+DAPO model across a configuration of NVIDIA L20 GPUs, leveraging the optimizations detailed in Section 3.5. This combination of model architecture and system-level enhancements reflects a design tailored to the practical constraints of large-scale deployment environments in real-world industrial scenarios. NES consistently achieved an average response time of under 450ms, delivering a seamless user experience while handling tens of thousands of daily calls. These results confirm the model’s effectiveness and reliability for industrial-grade applications.
6. Related Work
Transform-based Code Edits:
Several transformer-based approaches have been proposed to extend code generation solutions to code editing tasks (Li et al., 2023; Gupta et al., 2023; Liu et al., 2024a; Wei et al., 2024; Zhang et al., 2022a). CODIT (Chakraborty et al., 2018) utilizes AST-based operations for code edits but faces challenges with high computation and limited semantic understanding. SARGAM (Liu et al., 2023) introduces a three-stage framework (Search-Generate-Modify) combining retrieval, generation, and refinement for improved accuracy and efficiency. CodeT5 (Wang et al., 2021) is a versatile encoder-decoder model pre-trained for various code tasks. Although these methods differ in how they represent edits within deep learning models, they commonly frame the edit generation problem as a translation task, which often limits their ability to deliver an improved user experience.
General LLM-Powered Code Edits:
In recent years, general LLMs have made significant progress in domains such as code generation (Chen et al., 2021; Zeng et al., 2025; Di et al., 2023; Jiang et al., 2024; Tian and Chen, 2023; AI, 2025), code completion (Liu et al., 2017; Alon et al., 2019; Feng et al., 2020; Liu et al., 2020; Chakraborty and Ray, 2021; Lu et al., 2022; Liang et al., 2024; Naznin and Mondal, 2023; Wang et al., 2024a; Wan et al., 2024; Wang et al., 2024b), and other code-related tasks (Liu et al., 2024b). Codex (Chen et al., 2021) has demonstrated the ability to generate entire functions and applications from complex natural language prompts. CodeFuse-13B (Di et al., 2023) is an open-sourced pre-trained Language Model (LLM) with 13 billion parameters designed for code-related tasks. The latest open-source model, Kimi K2 (AI, 2025), has achieved state-of-the-art (SOTA) results among open-source models on both the SWE-bench Verified (Aleithan et al., 2024) and ACeBench  (Chen et al., 2025), demonstrating its leading capabilities in code generation task. LLM-powered tools are well-suited for general-purpose coding tasks but often lack optimization and specialized techniques for code editing tasks, which demand stricter requirements such as low latency, high accuracy, and efficiency.
Specialized Code Edit:
The advent of LLMs has spurred research in code editing, leading to the development of numerous approaches tailored specifically for code editing tasks (Li et al., 2023; Gupta et al., 2023; Liu et al., 2024a; Wei et al., 2024; Jin et al., 2025). InstructCoder (Li et al., 2023) introduces a dataset of 114,000 instruction-based triplets for LLM tuning, enhancing code editing accuracy for open-source models. However, it requires precise developer input, adding cognitive overhead. GRACE (Gupta et al., 2023) conditions LLMs on prior edits, significantly improving performance for edits distant from the current focus by capturing implicit developer intent. CoEdPilot (Liu et al., 2024a) combines historical edits, project context, and feedback to enhance code edit recommendations. It improves practicality but depends heavily on developer input and has high computational costs. Coeditor (Wei et al., 2024) uses repository-level changes and edit history to boost performance in single- and multi-turn editing. However, it may miss fine-grained edit patterns by focusing on repository diffs. Unlike these models that rely on developer prompts, NES takes a proactive approach by predicting edit locations and content. This enables a seamless, instruction-free workflow, fostering predictive collaboration.
7. Conclusion and Future Work
This paper presents NES, an LLM-powered framework for instruction-free and low-latency code editing assistance in IDEs. Unlike traditional approaches requiring explicit natural language instructions, NES leverages historical editing patterns to deliver efficient, context-aware support. It features two key components: the NES-Location Model, which predicts the next likely code modification location, and the NES-Edit Model, which generates contextually relevant code change suggestions based on editing history.
We evaluate NES using extensive real-world datasets from both open-source and industrial projects, highlighting its practical effectiveness. Experimental results show that the NES-Location Model achieves 75.6% accuracy in predicting the next edit location, while NES achieves 75.6% and 81.6% accuracy in two tasks of predicting next edit locations, alongside 91.36% ES and 27.7% EMR for intent-aligned edits, outperforming SOTA models. Also, our open-sourced SFT and DAPO datasets have been demonstrated to enhance the performance of open-source CodeLLMs.
Currently, NES operates within a single file, leveraging granular editing history in a localized context. Future work will enhance NES to handle multi-file code dependencies, enabling neural edit suggestions across interconnected files instead of single-file contexts.

### v3

@@
-An Efficient and Adaptive Next Edit Suggestion Framework with Zero Human Instructions in IDEs
+NES: An Instruction-Free, Low-Latency Next Edit Suggestion Framework Powered by Learned Historical Editing Trajectories
 Abstract.
-Code editing, including modifying, refactoring, and maintaining existing code, is the most frequent task in software development and has garnered significant attention from AI-powered tools.
-However, existing solutions that translate explicit natural language instructions into code edits face critical limitations, such as heavy reliance on human instruction input and high latency, which hinder their effective integration into a developer’s workflow.
-We observe that developers’ habitual behaviors and coding objectives are often reflected in their historical editing patterns, making this data key to addressing existing limitations.
-To leverage these insights, we propose NES (Next Edit Suggestion), an LLM-driven code editing framework that delivers an instruction-free and low-latency experience.
-Built on a dual-model architecture and trained with our high-quality SFT and DAPO datasets, NES enhances productivity by understanding developer intent while optimizing inference to minimize latency.
-NES is a scalable, industry-ready solution with a continuous Tab key interaction workflow, seamlessly adopted by a FinTech company with over 20,000 developers.
-Evaluations on real-world datasets show NES achieves 75.6% and 81.6% accuracy in two tasks of predicting next edit locations, alongside 91.36% ES and 27.7% EMR for intent-aligned edits, outperforming SOTA models.
-Our open-sourced SFT and DAPO datasets have been demonstrated to enhance the performance of open-source CodeLLMs.
-The demonstration of NES is available at https://youtu.be/yGoyYOe6fbY.
-Code Edit, Reinforcement Learning, Large Language Models, Software Engineering
-ccs: Software and its engineering
+Code editing is a frequent yet cognitively demanding task in software development.
+Existing AI-powered tools often disrupt developer flow by requiring explicit natural language instructions and suffer from high latency, limiting real-world usability.
+We present NES (Next Edit Suggestion), an instruction-free, low-latency code editing framework that leverages learned historical editing trajectories to implicitly capture developers’ goals and coding habits.
+NES features a dual-model architecture: one model predicts the next edit location, and the other generates the precise code change—both without any user instruction.
+Trained on our open-sourced SFT and DAPO datasets 2, NES achieves state-of-the-art performance (75.6% location accuracy, 27.7% exact match rate) while delivering suggestions in under 250ms.
+Deployed at Ant Group, NES serves over 20,000 developers through a seamless Tab-key interaction, achieving effective acceptance rates of 51.55% for location predictions and 43.44% for edits—demonstrating its practical impact in real-world development workflows.
+Code Edit, Reinforcement Learning (RL), Edit Location Prediction
+journalyear: 2026copyright: ccconference: 34th ACM Joint European Software Engineering Conference and Symposium on the Foundations of Software Engineering; July 05–09, 2026; Montreal, QC, Canadabooktitle: 34th ACM Joint European Software Engineering Conference and Symposium on the Foundations of Software Engineering (FSE Companion ’26), July 05–09, 2026, Montreal, QC, Canadadoi: 10.1145/3803437.3805244isbn: 979-8-4007-2636-1/2026/07ccs: Software and its engineering Software development techniquesccs: Computing methodologies Machine learning
 1. Introduction
-With Large Language Models (LLMs) demonstrating remarkable capabilities in various aspects of software development (Di et al., 2023; Yang et al., 2025; Wang et al., 2024b; Guo et al., 2024; Liang et al., 2024; Rozière et al., 2023; Ding et al., 2024; Codefuse et al., 2025), AI-powered development tools such as GitHub Copilot (Copilot, 2023), Cursor (Cursor, 2024) and Zeta (Zed, 2025) are revolutionizing modern software development workflows by leveraging LLMs for code editing tasks (Wang et al., 2021; Zhang et al., 2022b; Gupta et al., 2023; Wei et al., 2024; Liu et al., 2024a).
-Empirical studies further highlight that code editing, which includes activities such as modifying (Li et al., 2024; Guo et al., 2025), refactoring, and maintaining existing code, is one of the most frequent and critical activities throughout the lifecycle of a software project (Nguyen et al., 2013).
+With Large Language Models (LLMs) demonstrating remarkable capabilities in various aspects of software development (Di et al., 2023; Team, 2025b; Wang et al., 2024b; Guo et al., 2024; Liang et al., 2024; Rozière et al., 2023; Ding et al., 2024; Codefuse et al., 2025), AI-powered development tools such as GitHub Copilot (Copilot, 2023), Cursor (Cursor, 2024) and Zeta (Zed, 2025) are revolutionizing modern software development workflows by leveraging LLMs for code editing tasks (Wang et al., 2021; Zhang et al., 2022a; Gupta et al., 2023; Wei et al., 2024; Liu et al., 2024a).
+Empirical studies further highlight that code editing, which includes activities such as modifying (Li et al., 2023; Guo et al., 2025), refactoring, and maintaining existing code, is one of the most frequent and critical activities throughout the lifecycle of a software project (Nguyen et al., 2013).
+In particular, code completion task can be formally regarded as a restricted instance of code editing, constrained to insertions immediately following the cursor position within a partially written syntactic construct (e.g., statement, expression, or function body).
 The importance of efficient and accurate code editing cannot be overstated, as it directly impacts the productivity and quality of software development.
@@
 However, these approaches face critical limitations that hinder their effective integration into a developer’s workflow.
-On one hand, most existing methods rely heavily on human instruction input to maintain a coherent reference chain, disrupting programming flow and increasing cognitive load.
+On one hand, most existing methods, such as CodeT5 (Wang et al., 2021), CoditT5 (Zhang et al., 2022a), Coeditor (Wei et al., 2024), CoEdPilot (Liu et al., 2024a), EfficientEdit (Wang et al., 2025) and so on, rely heavily on human instruction input to maintain a coherent reference chain, disrupting programming flow and increasing cognitive load.
 Additionally, tools that fail to capture users’ habitual coding behavior often generate suggestions misaligned with their natural coding style.
 On the other hand, code editing tasks are inherently time-sensitive, with users expecting feedback within 1s(Card, 2018).
-Existing tools, however, rely on computationally intensive processes (e.g., search, natural language understanding), creating a tradeoff between low-latency and precise suggestions.
+But existing tools (Team, 2025b, a; Codefuse et al., 2025; Anthropic, 2024; Guo et al., 2024) rely on computationally intensive processes (e.g., search, text understanding, reasoning), creating a tradeoff between low-latency and precise suggestions.
 This significant latency undermines real-time interaction and degrades user experience.
 Overcoming these limitations can undoubtedly enhance code editing in real-world development.
-However, it necessitates modifying the method of acquiring cues, which can pose significant challenges in accurately identifying the editing location and determining the content within the desired timeframe.
+However, it necessitates modifying the method of acquiring cues, which poses significant challenges in accurately identifying the editing location and determining the content within the desired timeframe.
 Our approach is based on an empirical observation: users’ habitual behaviors and coding objectives are often embedded within their historical editing patterns (e.g., repetitive refactoring actions or managing cross-file dependencies), which provide valuable contextual cues.
@@
 We propose NES (Next Edit Suggestion), an LLM-driven code editing framework designed for an instruction-free and low-latency (feel like zero-latency) experience.
-Comprising a synergistic dual-model architecture trained on the proposed Supervised Fine-Tuning (SFT) and Dynamic sAmpling Policy Optimization (DAPO) datasets, NES adapts to user habits, integrates seamlessly into workflows, and enhances productivity by implicitly learning developer intent.
+Comprising a synergistic dual-model architecture trained on the proposed Supervised Fine-Tuning (SFT) (Wei et al., 2022) and Dynamic sAmpling Policy Optimization (DAPO) (Yu et al., 2025) datasets, NES adapts to user habits, integrates seamlessly into workflows, and enhances productivity by implicitly learning developer intent.
   • NES-Location Model utilizes the developer’s historical editing patterns to predict the most probable next editing location, facilitating proactive and efficient navigation across the codebase.
-  • NES-Edit Model analyzes the developer’s editing history, including refactoring patterns and API call sequences, to generate personalized and precise code modifications for the current editing region.
-To improve user experience, NES leverages Prefix Caching (PC) (Zheng et al., 2023) and Speculative Decoding (SD) (Leviathan et al., 2022; Chen et al., 2023) techniques to achieve substantial inference speedups.
-NES provides a scalable, industry-ready solution through its continuous Tab key interaction workflow, enabling seamless adoption in large-scale environments.
-Deployed in a FinTech company with 20,000+ developers, it handles thousands of daily code changes effortlessly.
-By allowing developers to perform complex tasks like refactoring or API call chains through simple "TabTabTab " actions, NES eliminates manual instructions, reduces cognitive load, and minimizes context-switching, delivering an efficient and immersive coding experience.
-We evaluate NES using extensive real-world datasets from both open-source and industrial projects, highlighting its practical effectiveness.
-Experimental results show that the NES-Location Model achieves 75.7% accuracy in predicting the next edit location, while the NES-Edit Model delivers a 27.7% exact match rate (EMR) for generating intent-aligned code edits, outperforming several state-of-the-art (SOTA) models.
-Moreover, NES reduces inference time to under 450ms by optimizing model performance on Nvidia L20 GPUs.
-Finally, training on our proposed SFT and DAPO datasets results in notable improvements on Qwen2.5-Coder-7B and Seed-Coder-8B-Instrcut models.
+  • NES-Edit Model analyzes the developer’s editing history, including refactoring patterns and API call sequences, to generate personalized and precise code modifications for the current edit.
+To improve user experience, NES leverages Prefix Caching (PC) (Zheng et al., 2024) and Speculative Decoding (SD) (Leviathan et al., 2022; Chen et al., 2023) techniques to achieve substantial inference speedups.
+Proven at scale with over 20,000 developers at Ant Group, NES is an industry-ready solution that streamlines complex coding tasks.
+Its core innovation is a continuous workflow where developers use simple Tab key sequences (TabTabTab) to perform operations like refactoring.
+This unique interaction model eliminates manual steps, significantly reducing cognitive load and context-switching to create a highly efficient and immersive coding experience.
+We demonstrate the practical effectiveness of NES using both open-source and industrial datasets.
+Our model achieves SOTA performance, with the Location Model reaching 75.6% accuracy in predicting edit placement and the Edit Model delivering a 27.7% Exact Match Rate (EMR) for generating intent-aligned code.
+By optimizing for Nvidia L20 GPUs, we ensure efficiency with an inference time of under 250ms.
+Additionally, our proposed SFT and DAPO datasets significantly improve existing Qwen2.5-Coder-7B and Seed-Coder-8B-Instruct models.
 Our key contributions are as follows.
-  • We propose NES  [^1] NES mentioned in the paper, unless specifically noted otherwise, use the Qwen3-4B+SFT+DAPO model.
-See Section 5 for details., a dual-model, LLM-driven code editing framework that leverages historical editing patterns and operates in an instruction-free, low-latency paradigm, eliminating the need for explicit natural language human input.
-  • We build the high-quality SFT and DAPO datasets [^2] We have submitted the datasets to GitHub and will make them publicly available soon at https://anonymous.link.for.review,which offer significant improvements in performance.
-These datasets have been instrumental in achieving accuracy and similarity scores that are dozens of times higher on code editing tasks across a variety of models.
-  • We have successfully integrated and deployed NES on Nvidia L20 GPUs within a FinTech organization.
-Utilizing a comprehensive set of inference optimizations, NES has achieved low-latency performance, with response times under 450ms, thereby enhancing the user experience.
+  • We introduce NES  [^1] Unless otherwise specified, NES refers to the Qwen3-4B+SFT+DAPO model.
+See Section 5 for further details., a dual-model, LLM-powered code editing framework that leverages historical code edit patterns.
+Operating in an instruction-free, low-latency setting, it eliminates the need for explicit natural language input from users.
+  • We construct and open-source high-quality SFT and DAPO datasets [^2] A portion of the dataset is publicly available at https://huggingface.co/datasets/codefuse-ai/CodeFuse_codeedit.
+The remaining data will be released incrementally., which significantly boost performance.
+These datasets have played a critical role in achieving dramatic improvements—by orders of magnitude—in accuracy and similarity metrics across various code editing benchmarks.
+  • We have integrated and deployed NES across Nvidia L20 GPUs within Ant Group, serving 20,000+ developers.
+The implementation of various inference optimizations ensures that NES delivers real-time responses (under 250ms), culminating in a more fluid user experience with effective acceptance rates of 43.44% for code edits and 51.55% for location prediction.
 2. Motivating Example
@@
 The process involves manual navigation, repetitive edits, and significant cognitive effort to track changes.
-Although LLM-driven tools can assist with locating and making edits, instruction-based tools, such as CodeT5 (Wang et al., 2021), CoditT5 (Zhang et al., 2022b), Coeditor (Wei et al., 2024), and CoEdPilot (Liu et al., 2024a), disrupt the developer’s workflow by requiring precise commands and waiting for responses, introducing additional cognitive overhead.
+Although LLM-driven tools can assist with locating and making edits, instruction-based tools, such as CodeT5 (Wang et al., 2021), CoditT5 (Zhang et al., 2022a), Coeditor (Wei et al., 2024), and CoEdPilot (Liu et al., 2024a), disrupt the developer’s workflow by requiring precise commands and waiting for responses, introducing additional cognitive overhead.
 To address these challenges, an instruction-free tool is essential, which is where NES comes in.
@@
 Here, NES-Edit suggests context-specific values (e.g., aria-label="Save Changes") for the new property, with each step accepted via Tab.
-A tedious, error-prone manual process is streamlined into an efficient "TabTabTab " workflow.
+A tedious, error-prone manual process is streamlined into an efficient ”TabTabTab ” workflow.
 Powered by high-quality datasets, NES predicts both the location (NES-Location) and content (NES-Edit) of edits without requiring explicit instructions.
@@
 The framework’s functionality is initiated by the developer’s initial editing action, subsequently entering a dynamic interactive cycle:
-  • Next Location Suggestion: The NES-Location Model, leveraging real-time updated historical editing trajectories ①, predicts the developer’s potential next edit location.
-Subsequently, the framework presents this predictive suggestion to the developer, who then has the option to either rapidly navigate to the suggested position by pressing the Tab key or to disregard the suggestion ②.
-  • Code Edit Suggestion: Thereafter, the NES-Edit Model, integrating historical editing patterns with the current editing window’s context, generates real-time code editing suggestions ③.
-Developers have the flexibility to either accept these suggestions by pressing the Tab key or continue with independent edits ④, while also updating the edit history ⑤.
-  • Dynamic loop: Upon the developer’s adoption of an editing suggestion, the framework promptly initiates the next round of next location suggestion, establishing a continuous intelligent assistance loop.
-Ideally, developers can significantly enhance their coding efficiency and quality by consistently accepting NES framework’s suggestions by pressing the Tab key ⑥.
+  • Next Location Suggestion: The NES-Location Model leverages real-time editing trajectories ① to predict the next likely edit location.
+The developer can rapidly jump to this position by pressing the Tab key or simply disregard the suggestion ②.
+  • Code Edit Suggestion: Once at the location, the NES-Edit Model generates a code edit by integrating historical patterns with the current code context ③.
+This suggestion can be accepted with the Tab key, or the developer can continue editing manually ④, which in turn updates the edit history ⑤.
+  • Dynamic loop: Each accepted suggestion triggers the next prediction, creating a continuous assistance loop ⑥.
+This process transforms complex refactoring tasks into a fluid Tab sequence, significantly boosting developer productivity.
 To implement this framework, we focused on four key areas:
-  • Incremental Difference Detection: We developed a real-time code change monitor based on incremental computation.
+Incremental Difference Detection
+We developed a real-time code change monitor based on incremental computation.
 This framework accurately captures each developer editing operation, thereby constructing a comprehensive and detailed historical editing trajectory.
 These trajectory data serve as crucial inputs for subsequent model training and prediction processes.
-  • High-Quality Training Dataset Construction: Utilizing the collected authentic historical editing trajectories, we constructed a high-quality training dataset.
+High-Quality Training Dataset Construction
+Utilizing the collected authentic historical editing trajectories, we constructed a high-quality training dataset.
 This dataset encompasses not only sequences of editing operations but also contextual information for each edit.
-Therefore, this dataset can enabling the training of both the NES-Location Model and the NES-Edit Model with nuanced understanding of editing patterns and their surrounding code environments.
-  • Model Training: To endow NES with the capability for location and code editing suggestion, we designed a two-stage training methodology.
+Therefore, this dataset can enable the training of both the NES-Location Model and the NES-Edit Model with nuanced understanding of editing patterns and their surrounding code environments.
+Model Training
+To endow NES with the capability for location and code editing suggestion, we designed a two-stage training methodology.
 Initially, both the NES-Location and NES-Edit models undergo SFT (Ouyang et al., 2022) on our large-scale historical editing datasets, enabling them to learn the fundamental patterns of code modification.
-Subsequently, we employ Decoupled Clip and DAPO (Yu et al., 2025), a reinforcement learning technique using high-quality preference data, to further refine the models.
+Subsequently, we employ Decoupled Clip and DAPO (Yu et al., 2025), a reinforcement learning technique using high-quality preference data, to further refine the models [^3] DAPO is a reinforcement learning method that uses dynamic sampling of high-quality preferences to refine the policy, while Decoupled Clip stabilizes training by decoupling the clipping mechanism in policy updates..
 This second stage aligns their behavior more closely with real-world developer intent and utility, significantly boosting the practical value of NES.
-  • Model Inference Optimization: Acknowledging the constraints of inference resources in practical application environments and the stringent low-latency requirements inherent in real-time editing suggestion, we conducted in-depth optimization of the model inference process, and achieved real-time responsiveness to developer editing behaviors, ensuring the fluidity and practicality of the NES framework.
+Model Inference Optimization
+Acknowledging the constraints of inference resources in practical application environments and the stringent low-latency requirements inherent in real-time editing suggestion, we conducted in-depth optimization of the model inference process, and achieved real-time responsiveness to developer editing behaviors, ensuring the fluidity and practicality of the NES framework.
 We present the details of these critical components in the following sections, respectively.
 3.2. Incremental Difference Detection
-We developed an efficient incremental difference detection mechanism built upon our AI-powered IDE, designed to accurately capture real-world code editing trajectories of developers during real-world programming activities.
-As detailed in Algorithm  1, it operates via two steps: incremental difference calculation and difference merging, where "Overlap" denotes intersecting regions of two editing operations.
+We developed an efficient incremental difference detector built upon our AI-powered IDE, designed to accurately capture real-world code editing trajectories of developers during real-world programming activities.
+The algorithm operates in two primary steps: incremental difference calculation and difference merging, where ”Overlap” denotes intersecting regions of two editing operations.
 3.2.1. Incremental Difference Calculation
@@
 To address this, we implemented instant merging of overlapping differences, using developer-edited code blocks as units to create cohesive recommendations aligned with real-world editing behaviors.
+Table 1. An example comparison between the diff formats of difflib and NES highlights that NES’s format offers higher information density with precise line numbers.
 difflib format NES diff format
-@@ -1,3 +1,3 @@ -def Hello() +def GoodBye() print("Say") - print("Hello") + print("GoodBye") 1-| def Hello() 1+| def GoodBye() 2 | print("Say") 3-| print("Hello") 3+| print("GoodBye")
-Table 1. An example comparison between the diff formats of difflib and NES highlights that NES’s format offers higher information density with precise line numbers.
-Algorithm 1 Incremental Difference Detecting
-A stream of edit events $E_{stream}$, where each event $e=(C_{pre},C_{post})\in E_{stream}$ encompasses the pre- and post-modification content of code snippets.
-A list of merged historical edits =$\mathcal{H}=\{\Delta_{1},\Delta_{2},\dots\}$.
-Initialize $\mathcal{H}\leftarrow\emptyset$ and $\Delta_{active}\leftarrow\text{null}$.
-for each incoming edit event $(C_{pre},C_{post})$ in $E_{stream}$ do
-  $\Delta_{new}\leftarrow\text{Diff}(C_{pre},C_{post})$
-  if $\Delta_{active}$ is null then
-    $\Delta_{active}\leftarrow\Delta_{new}$
-  else if $\text{Overlap}(\Delta_{active},\Delta_{new})$ then
-    $\Delta_{active}\leftarrow\text{Merge}(\Delta_{active},\Delta_{new})$
-  else
-    Add $\Delta_{active}$ to $\mathcal{H}$
-    $\Delta_{active}\leftarrow\Delta_{new}$
-  end if
-end for
-After the stream ends, add the last non-null $\Delta_{active}$ to $\mathcal{H}$.
-return the list of historical trajectories $\mathcal{H}$.
+@@ -1,3 +1,3 @@ -def Hello() +def GoodBye() print(”Say”) - print(”Hello”) + print(”GoodBye”) 1-— def Hello() 1+— def GoodBye() 2 — print(”Say”) 3-— print(”Hello”) 3+— print(”GoodBye”)
 Figure 4. Training dataset collection process
@@
 To mitigate this, we introduce a crucial data refinement stage.
-We employ an LLM as a relevance filter to analyze the semantic and causal relationship between the historical edits ($H_{T}$) and the current edit ($L_{\text{gt}}$, $E_{\text{gt}}$).
+We employ an LLM, like Qwen, as a relevance filter to analyze the semantic and causal relationship between the historical edits ($H_{T}$) and the current edit ($L_{\text{gt}}$, $E_{\text{gt}}$).
 Based on the LLM’s assessment, each instance is processed as follows:
@@
 This crucial step teaches the model to refrain from making a suggestion when the developer’s intent is not inferable from the available history.
-This filtering and relabeling process transforms the preliminary training datacset into a final, optimized training set.
+This filtering and relabeling process transforms the preliminary training dataset into a final, optimized training set.
 This refined dataset not only enhances the quality of modification-oriented samples but, critically, also incorporates explicit negative samples.
@@
 A reward of 1.0 is granted for an exact match, strongly encouraging the model to pursue perfect accuracy.
-If the generated edit is not an exact match but is substantially similar to the ground truth (ES > 0.5), a proportional reward is assigned.
+If the generated edit is not an exact match but is substantially similar to the ground truth (ES ¿ 0.5), a proportional reward is assigned.
 The scaling factor of 0.5 ensures this partial credit is strictly less than the reward for a perfect solution, while still guiding the model to generate semantically relevant code.
@@
 4. Evaluation
-We evaluate NES by addressing the following research questions:
+We evaluate NES by addressing the following questions:
   • RQ1 Next Location Prediction: To what extent can NES accurately predict the position of the next edit?
@@
 For RQ1, which evaluates the prediction of the next editing position, the primary metric used is Accuracy (ACC), measuring whether the predicted edit location matches the ground truth.
-For RQ2, which evaluates code-editing performance, the three key metrics are: Edit Similarity (ES) for semantic alignment between the generated and target edits; Exact Match Rate (EMR) for syntactic precision of the generated edits for the modification task; and Accuracy (ACC) for correctly predict preservation.
+For RQ2, which evaluates code-editing performance, the three key metrics are: Edit Similarity (ES) (Song et al., 2024) for semantic alignment between the generated and target edits; Exact Match Rate (EMR) for syntactic precision of the generated edits for the modification task; and Accuracy (ACC) for correctly predict preservation.
 For RQ3, which investigates model inference speed, the performance metrics are input and output tokens/s speed.
 4.1.3. Baseline Models
-We compare NES to the following models.
-  • Qwen (Yang et al., 2025), developed by Alibaba Cloud, is a renowned LLM, while Qwen-Coder (Yang et al., 2024) specializes in programming and software development tasks.
-  • Claude (Anthropic, [n. d.]), created by Anthropic, has strong capabilities in coding and can assist with a wide range of programming tasks, including writing, debugging, and explaining code.
+We compare NES to the models.
+  • Qwen (Team, 2025b), developed by Alibaba Cloud, is a renowned LLM, while Qwen-Coder (Team, 2024) specializes in programming and software development tasks.
+  • Claude (Anthropic, 2024), created by Anthropic, has strong capabilities in coding and can assist with a wide range of programming tasks, including writing, debugging, and explaining code.
   • Seed-Coder (Seed et al., 2025), recently released by ByteDance, is a lightweight yet powerful open-source code-focused large language model (LLM) that delivers strong performance across a variety of coding benchmarks.
@@
 Zeta intuitively generates context-aware code or text modifications based solely on the input provided.
-We categorize these models into two groups based on their size.
+  • CoEdPilot (Liu et al., 2024a) jointly trains a localization LLM (to pinpoint edits) and a generation LLM.
+However, the reliance on explicit human instructions constrains their overall editing efficiency.
 The models with relatively smaller sizes include Qwen3-4B, Qwen2.5-Coder-7B, and Seed-Coder-8B-Instruct.
@@
 Training large models is resource-intensive and consumes significant computational resources.
-We did not select instruction-based tools such as CodeT5 (Wang et al., 2021), CoditT5 (Zhang et al., 2022b), Coeditor (Wei et al., 2024), and CoEdPilot (Liu et al., 2024a) as baselines, as these tools rely on explicit human instructions to operate.
-Since our evaluation environment and benchmarks do not account for such input, including these tools would lead to an unfair comparison.
-However, based on feedback from real-world developers, our approach provides an editing tool that is at least as effective, if not superior, in delivering practical assistance to users.
+We select CoEdPilot (Liu et al., 2024a) as the baseline from the family of explicit human-instruction-based code editing tools, alongside models like CodeT5 (Wang et al., 2021), CoditT5 (Zhang et al., 2022a), Coeditor (Wei et al., 2024), EfficientEdit (Wang et al., 2025) and so on.
+As our evaluation environment does not natively support CoEdPilot’s required human instruction input format, we left the corresponding input field blank during testing.
 4.1.4. Settings
@@
 4.2. Results
+Table 2.
+Accuracy comparison for location tasks (-do and -keep) across models.
+“model+SFT” and “model+SFT+DAPO” indicate training on the base model with the NES SFT dataset alone and on both the SFT and DAPO datasets.
 Model TypeScript TypeScriptReact Java Python Average
@@
 Claude4 65.2% 8.0% 65.0% 10.0% 60.0% 15.6% 77.6% 4.7% 67.0% 9.6%
+CoEdPilot 63.2% 4.4% 62.6% 3.0% 12.8% 6.4% 17.6% 27.4% 31.6% 8.2%
 Qwen2.5-Coder-7B+SFT 62.4% 94.6% 69.3% 97.0% 51.3% 95.3% 68.4% 89.3% 62.8% 94.0%
@@
 Qwen3-4B+SFT+DAPO 76.5% 82.3% 77.3% 86.2% 71.8% 80.1% 76.9% 77.8% 75.6% 81.6%
-Table 2.
-Accuracy comparison for location tasks (-do and -keep) across models.
-“model+SFT” and “model+SFT+DAPO” indicate training on the base model with the NES SFT dataset alone and on both the SFT and DAPO datasets.
+* Training large models on more data was infeasible because of the necessary low-latency performance and restrictive training resource constraints.
+Zeta was not evaluated because it lacks navigation functionality.
+CoEdPilot was evaluated without human instructions.
+Table 3. Comparison of location prediction performance with varying max edit history lengths on the Qwen3-4B+SFT model
 Max Edit History TypeScript TypeScriptReact Java Python Average
@@
 9 67.6% 82.2% 71.4% 86.2% 60.0% 80.1% 72.1% 75.7% 67.8% 81.0%
-Table 3. Comparison of location prediction performance with varying maximum edit history lengths on the Qwen3-4B+SFT model
 4.2.1. RQ1 NES-Location:
@@
 The terms “model+SFT” and “model+SFT+DAPO” indicate training on the base model with the NES SFT dataset alone and on both the NES SFT and DAPO datasets, respectively.
-Impact of SFT: Before post-training, the best base model, Claude 4, achieves 67.0% and 9.6% average accuracy on the -do and -keep tasks, respectively.
+Before post-training, the best base model, Claude4, achieves 67.0% and 9.6% average accuracy on the -do and -keep tasks, respectively.
 Models with small size, such as Qwen2.5-Coder-7B, perform poorly, with only 10.1% and 3.7% on the -do and -keep tasks, respectively.
 NES SFT dataset significantly improves models with small size.
-For -do, Qwen2.5-Coder-7B jumps from 10.1% to 62.8%, Seed-Coder-8B-Instruc from 9.3% to 72.6%, and Qwen3-4B from 6.6% to 72.6%.
-For -keep, Qwen2.5-Coder-7B rises from 3.7% to 94.0%, Seed-Coder-8B-Instruc from 2.9% to 94.1%, and Qwen3-4B from 1.9% to 89.4%.
-Due to the limits of low-latency requirements and training resource constraints, large models were not trained on the SFT dataset.
-Additionally, Zeta does not provide navigation functionality, so its results are excluded from Table 2.
-RQ1 Answer about SFT: The NES SFT dataset significantly boosts model performance in predicting the next editing position, improving Qwen3-4B’s accuracy from 6.6% to 72.6% on the -do task and from 1.9% to 89.4% on the -keep task.
-Impact of DAPO: Models with smaller sizes trained on the DAPO dataset achieve slight improvements in average navigation prediction (-do) accuracy: Qwen2.5-Coder-7B improves from 62.8% to 65.6%, Seed-Coder-8B-Instruc increases from 72.6% to 77.7%, and Qwen3-4B rises from 72.6% to 75.6%.
-However, on the -keep task, Qwen2.5-Coder-7B declines from 94.0% to 86.5%, Seed-Coder-8B-Instruct declines from 94.1% to 85.0%, and Qwen3-4B drops from 89.4% to 81.6%.
+For -do, Qwen2.5-Coder-7B jumps from 10.1% to 62.8%, Seed-Coder-8B-Instruct from 9.3% to 72.6%, and Qwen3-4B from 6.6% to 72.6%.
+For -keep, Qwen2.5-Coder-7B rises from 3.7% to 94.0%, Seed-Coder-8B-Instruct from 2.9% to 94.1%, and Qwen3-4B from 1.9% to 89.4%.
+Limits.
+Due to constraints on low-latency requirements and training resources, large models were not trained on the SFT or DAPO dataset.
+Furthermore, since Zeta lacks navigation functionality, its results are excluded from Table 2.
+Finally, CoEdPilot’s reliance on explicit human inputs explains why its performance did not surpass that of other tools.
+A similar result was observed in the subsequent evaluation.
+Impact of SFT: The NES SFT dataset significantly boosts model performance in predicting the next editing position, improving Qwen3-4B’s accuracy from 6.6% to 72.6% on the -do task and from 1.9% to 89.4% on the -keep task.
+Training smaller models on the DAPO dataset results in a slight but notable improvement in average navigation prediction (-do) accuracy: Qwen2.5-Coder-7B improves from 62.8% to 65.6%, Seed-Coder-8B-Instruct increases from 72.6% to 77.7%, and Qwen3-4B rises from 72.6% to 75.6%.
+But, on the -keep task, Qwen2.5-Coder-7B declines from 94.0% to 86.5%, Seed-Coder-8B-Instruct declines from 94.1% to 85.0%, and Qwen3-4B drops from 89.4% to 81.6%.
 The DAPO-trained model tends to suggest location changes more frequently.
-Because changing is more frequently in the real-world development practices, We accept the minor drop in -keep performance in favor of improved -do accuracy.
-The NES-Location Model aims to streamline navigation, so we’ve adopted DAPO, prioritizing the higher success rate in predicting necessary jumps over avoiding suggestions, as it better supports developers’ workflow.
-RQ1 Answer about DAPO: The DAPO-trained model tends to suggest location changes more frequently, leading to a slight drop in -keep task accuracy.
-However, we have decided to adopt the DAPO-trained NES-Location model, which demonstrates better performance on the -do task, as location changes are more prevalent in real-world development practices.
-Impact of Edit History Length: Table 3 demonstrates the impact of historical context length (Max Edit History) on Qwen3-4B+SFT.
+This is because changing is more frequent in the real-world development practices.
+We accept the minor drop in -keep performance in favor of improved -do accuracy.
+DAPO was chosen because predicting necessary jumps accurately matters more than avoiding suggestions in developers’ workflow.
+Impact of DAPO: The DAPO-trained model tends to suggest location changes more frequently, leading to a slight drop in -keep task accuracy.
+However, the NES-Location model was chosen for its superior -do task performance, as location changes are more prevalent in practice.
+Table 3 demonstrates the impact of historical context length (Max Edit History) on Qwen3-4B+SFT.
 A history length of 3 achieves the best balance, with 72.6% -do average accuracy and 89.4% -keep average accuracy.
 Short histories (e.g., 1) lack sufficient context for complex intents, while longer histories (e.g., 5+) introduce irrelevant noise, hindering accuracy.
-RQ1 Answer about Edit History Length: Short histories may lack context, while long histories can introduce noise, both hindering accuracy. The results show that a history length of 3 is optimal.
+Impact of Edit History Length: Short histories may lack context, while long histories can introduce noise, both hindering accuracy. The results show that a history length of 3 is optimal.
+Table 4. Similarity comparison for modification tasks (-do). “model+SFT” and “model+SFT+DAPO” indicate training on the base model with the NES SFT dataset alone and on both the SFT and DAPO datasets.
 Model TypeScript TypeScriptReact Java Python Average
@@
 Qwen2.5-Coder-7B 72.17/12.4% 69.52/6.4% 64.22/4.6% 72.52/8.8% 69.61/8.1%
-Seed-Coder-8B-Instrcut 71.30/14.9% 69.35/7.0% 69.91/8.8% 70.06/9.8% 70.16/10.1%
+Seed-Coder-8B-Instruct 71.30/14.9% 69.35/7.0% 69.91/8.8% 70.06/9.8% 70.16/10.1%
 Qwen3-4B 76.77/6.4% 80.46/13.5% 73.32/6.6% 77.23/7.0% 76.94/8.5%
@@
 Claude4 85.37/25.7% 85.56/19.6% 83.40/22.4% 83.38/24.8% 84.43/23.1%
+CoEdPilot 72.34/4.3% 77.73/1.6% 72.06/0.8% 60.88/1.8% 70.75/2.1%
 Qwen2.5-Coder-7B+SFT 92.49/28.8% 92.63/28.8% 89.17/19.6% 89.61/22.6% 90.98/25.0%
-Seed-Coder-8B-Instrcut+SFT 92.29/30.2% 92.85/34.0% 89.55/20.4% 89.54/22.0% 91.06/26.7%
+Seed-Coder-8B-Instruct+SFT 92.29/30.2% 92.85/34.0% 89.55/20.4% 89.54/22.0% 91.06/26.7%
 Qwen3-4B+SFT 92.03/27.0% 90.69/29.6% 89.29/20.2% 89.30/21.6% 90.33/24.6%
 Qwen2.5-Coder-7B+SFT+DAPO 91.82/32.4% 92.67/34.4% 89.99/21.0% 90.03/25.4% 91.13/28.3%
-Seed-Coder-8B-Instrcut+SFT+DAPO 91.90/31.4% 93.21/34.2% 90.15/23.6% 90.08/26.0% 91.34/28.8%
+Seed-Coder-8B-Instruct+SFT+DAPO 91.90/31.4% 93.21/34.2% 90.15/23.6% 90.08/26.0% 91.34/28.8%
 Qwen3-4B+SFT+DAPO 92.77/30.6% 93.28/34.2% 89.97/21.4% 89.41/24.6% 91.36/27.7%
-Table 4. Similarity comparison for modification tasks (-do). “model+SFT” and “model+SFT+DAPO” indicate training on the base model with the NES SFT dataset alone and on both the SFT and DAPO datasets.
+Table 5. Accuracy comparison for preservation tasks (-keep). “model+SFT” and “model+SFT+DAPO” indicate training on the base model with the NES SFT dataset alone and on both the SFT and DAPO datasets.
 Model TypeScript TypeScriptReact Java Python Average
 Qwen2.5-Coder-7B 11.6% 8.0% 4.6% 8.8% 8.3%
-Seed-Coder-8B-Instrcut 2.6% 3.2% 2.0% 3.8% 2.9%
+Seed-Coder-8B-Instruct 2.6% 3.2% 2.0% 3.8% 2.9%
 Qwen3-4B 35.2% 37.4% 28.8% 44.8% 36.6%
@@
 Claude4 15.6% 15.2% 11.0% 17.0% 14.7%
+CoEdPilot 11.2% 4.6% 7.0% 23.2% 11.5%
 Qwen2.5-Coder-7B+SFT 94.4% 90.2% 87.4% 92.0% 91.0%
-Seed-Coder-8B-Instrcut+SFT 92.2% 86.0% 88.6% 89.8% 89.2%
+Seed-Coder-8B-Instruct+SFT 92.2% 86.0% 88.6% 89.8% 89.2%
 Qwen3-4B+SFT 92.4% 84.8% 88.0% 88.8% 88.5%
 Qwen2.5-Coder-7B+SFT+DAPO 84.8% 85.2% 80.8% 86.2% 84.3%
-Seed-Coder-8B-Instrcut+SFT+DAPO 85.2% 80.0% 78.2% 83.4% 81.7%
+Seed-Coder-8B-Instruct+SFT+DAPO 85.2% 80.0% 78.2% 83.4% 81.7%
 Qwen3-4B+SFT+DAPO 86.2% 88.8% 82.0% 88.2% 86.3%
-Table 5. Accuracy comparison for preservation tasks (-keep). “model+SFT” and “model+SFT+DAPO” indicate training on the base model with the NES SFT dataset alone and on both the SFT and DAPO datasets.
+Table 6. Similarity comparison with different ratios of unmodified data
 Ratio of TypeScript TypeScriptReact Java Python Average
@@
 60% 91.89/21.4% 98.6% 91.01/20.0% 96.0% 88.73/12.2% 95.4% 88.92/15.6% 95.8% 90.14/17.3% 96.5%
-Table 6. Similarity comparison with different ratios of unmodified data
+Table 7. Comparison of code edit generation with varying maximum edit history lengths on the Qwen3-4B+SFT model
 Max TypeScript TypeScriptReact Java Python Average
@@
 9 91.61/29.2% 93.0% 91.07/23.0% 87.8% 89.06/16.4% 90.2% 89.05/20.4% 92.2% 90.20/22.3% 90.8%
-Table 7. Comparison of code edit generation with varying maximum edit history lengths on the Qwen3-4B+SFT model
 4.2.2. RQ2 Code Edit Generation:
 How effectively can NES generate precise edit suggestions?
-Impact of SFT: Table 4 and Table 5 present the performance of the NES-Edit Model across four programming languages: Java, Python, TypeScript, and TypeScript React.
+Table 4 and Table 5 present the performance of the NES-Edit Model across four programming languages: Java, Python, TypeScript, and TypeScript React.
 In the -do task, Zeta achieves the highest ES score of 85.94%, while Claude4 records the best EMR score of 23.1%.
@@
 After SFT fine-tuning, smaller models show significant improvements in accuracy, with Qwen2.5-Coder-7B+SFT reaching 91.0%, Seed-Coder-8B-Instruct+SFT achieving 89.2%, and Qwen3-4B+SFT attaining 88.5%.
-RQ2 Answer about SFT: The NES SFT dataset greatly improves model performance in code edit generation.
-Impact of DAPO: Smaller models trained on the DAPO dataset achieve moderate improvements in average code edit generation (-do) similarity scores.
+Impact of SFT: The NES SFT dataset greatly improves model performance in code edit generation.
+Smaller models trained on the DAPO dataset achieve moderate improvements in average code edit generation (-do) similarity scores.
 For example, Qwen2.5-Coder-7B improves from an ES of 90.98% to 91.13% and an EMR of 25.0% to 28.3%; Seed-Coder-8B-Instruct increases from an ES of 91.06% to 91.34% and an EMR of 26.7% to 28.8%; and Qwen3-4B improves from an ES of 90.33% to 91.36% and an EMR of 24.6% to 27.7%.
-For the -keep task, Qwen2.5-Coder-7B experiences a decline in accuracy from 91.0% to 84.3%, Seed-Coder-8B-Instruct sees an decrease from 89.2% to 81.7%, and Qwen3-4B drops from 88.5% to 86.3%.
-The slight decrease in performance on the -keep task can be attributed to the same reason as with NES-Location: the DAPO-trained model tends to suggest location changes more frequently.
-Nevertheless, we deploy the DAPO-trained model in real-world scenarios due to its alignment with high-frequency practical use cases.
-RQ2 Answer about DAPO: The DAPO-trained model demonstrates improved similarity scores for the -do task, yet shows a minor drop in accuracy for the -keep task.
-However, we have chosen to adopt the model trained with DAPO, as it better aligns with the high-frequency practices observed in real-world software development.
-Impact of Unmodified Data: We investigated the impact of varying proportions of unmodified data in the training set, analyzing ratios from 10% to 60% (Table 6).
+For the -keep task, Qwen2.5-Coder-7B experiences a decline in accuracy from 91.0% to 84.3%, Seed-Coder-8B-Instruct drops from 89.2% to 81.7%, and Qwen3-4B drops from 88.5% to 86.3%.
+The slight -keep drop is due to the DAPO-trained model suggesting changes more frequently, but we deploy it for its alignment with high-frequency practical use cases.
+Impact of DAPO: The DAPO-trained model demonstrates improved similarity scores for the -do task, yet shows a minor drop in accuracy for the -keep task.
+However, we have chosen to adopt the model trained with DAPO, as it better aligns with the high-frequency practices observed in real-world development.
+We investigated the impact of varying proportions of unmodified data in the training set, analyzing ratios from 10% to 60% (Table 6).
 The results show that higher proportions of unmodified data improve performance on the -keep test sets, while leading to a decline in performance on the -do test sets.
@@
 As a result, we selected the 20% unmodified data ratio as the standard configuration for all subsequent experiments.
-RQ2 Answer about Unmodified Data: As a result, we selected the 20% unmodified data ratio as the standard configuration for all subsequent experiments.
-Impact of Edit History Length: Table 7 illustrates the effect of varying the length of historical context (Max Edit History) on the performance of Qwen3-4B+SFT.
-Optimal results are achieved with a history length of 3, yielding a -do average similarity score of ES 90.33% and EMR 24.6%.
-Conversely, a history length of 7 results in the highest -keep average accuracy of 88.5%.
-This suggests that generating better code edits may require a more extensive historical trajectory.
-RQ2 Answer about Edit History Length: Generating better code edits may require a more extensive historical trajectory.
+Impact of Unmodified Data: We selected the 20% unmodified data ratio as the standard configuration for all subsequent experiments.
+Table 7 shows that a history length of 3 is optimal for the -do task (ES 90.33%, EMR 24.6%), while a length of 7 yields the best -keep accuracy (91.7%), suggesting better code edits may require longer historical trajectories.
+Impact of Edit History Length: Generating better code edits may require a more extensive historical trajectory.
 4.2.3. RQ3 Model Performance:
 Does NES demonstrate sufficient computational efficiency to provide low-latency suggestions in interactive development environments?
-Table 8 presents the results of our inference speed analysis.
-The data reveals that both Prefix Caching (PC) and Speculative Decoding (SD) significantly enhance inference speed, with Speculative Decoding providing the largest improvement.
-This aligns with our initial hypothesis, confirming the suitability of Speculative Decoding for our specific use case.
-When compared to Zeta, a quantized 7B model with the same inference optimizations, NES achieves superior input throughput, while maintaining equivalent output throughput.
-In our online environment, NES ultimately attained an average inference latency of 450ms, effectively reaching real-time processing capabilities and successfully meeting our performance objectives.
+Table 8 shows that Prefix Caching (PC) (Zheng et al., 2024), for reusing prompt computations, and Speculative Decoding (SD) (Chen et al., 2023; Leviathan et al., 2022), for faster token generation, significantly boost performance, with SD showing the largest gains.
+Evaluated on vLLM, NES demonstrates superior input throughput over Zeta, while a latency comparison with CoEdPilot was deemed infeasible, as its non-vLLM-based architecture would require a costly re-implementation for an equitable benchmark.
+These optimizations enable NES to achieve an average latency of 250ms in live deployment.
+Table 8. Comparison of input and output token processing speeds for different model configurations.
 Model configurations input tokens/s output tokens/s
@@
 Zeta+vLLM+PC+SD 5400 125
-Table 8. Comparison of input and output token processing speeds for different model configurations
-RQ3 Answer: The Speculative Decoding (SD) optimization delivers the most significant improvement in inference input/output throughput, enabling NES to achieve an average inference latency of 450ms.
-5. Design of Industrial Tools
-In transitioning the NES framework from a research prototype to a scalable, industrial-grade solution, we faced the challenge of balancing model performance, inference latency, and deployment cost.
-To meet the stringent responsiveness demands of real-time code editing, we chose the Qwen3-4B model as the core engine for NES, post-train it with our SFT and DAPO datasets (see Section 3.3) for production deployment.
+* CoEdPilot’s non-vLLM architecture is incompatible with PC and SD optimizations.
+Answer to RQ3: The SD optimization delivers the most significant improvement in inference input/output throughput, enabling NES to achieve an average inference latency of 250ms.
+5. Industrial Deployment
+5.1. Design Rationale
+In transitioning the NES framework to a scalable, industrial-grade solution, we faced the challenge of balancing model performance, inference latency, and deployment cost.
+To meet the stringent responsiveness demands of real-time code editing, we chose the Qwen3-4B model as the core engine, post-train it with our SFT and DAPO datasets (see Section 3.3) for production deployment.
 This decision was based on several key considerations:
   • Trade-off between Model Size and Latency: Larger scale models, like 8B-parameter ones, have high computational needs and slower inference, making them unsuitable for real-time code editing’s low-latency requirements.
-The Qwen3-4B model, being smaller, offers faster inference, aligning better with our goal of achieving a "zero-latency" user experience in NES.
+The Qwen3-4B model, being smaller, offers faster inference, aligning better with our goal of achieving a ”zero-latency” user experience in NES.
   • Excellent Cost-Performance Ratio after DAPO Training: As detailed in Section 4, our two-stage training process greatly improved Qwen3-4B’s performance.
@@
   • Feasibility for Large-Scale Deployment: In industrial applications, ensuring service stability and scalability is critical.
-We deployed the Qwen3-4B+SFT+DAPO model across a configuration of NVIDIA L20 GPUs, leveraging the optimizations detailed in Section 3.5.
-This combination of model architecture and system-level enhancements reflects a design tailored to the practical constraints of large-scale deployment environments in real-world industrial scenarios.
-NES consistently achieved an average response time of under 450ms, delivering a seamless user experience while handling tens of thousands of daily calls.
-These results confirm the model’s effectiveness and reliability for industrial-grade applications.
+We deployed the Qwen3-4B+SFT+DAPO model across a configuration of NVIDIA L20 GPUs with the optimizations in Section 3.5, tailored for large-scale deployment.
+NES consistently achieved an average response time of under 250ms, delivering a seamless user experience while handling tens of thousands of daily calls.
+These results confirm the model’s effectiveness and reliability for industrial use.
+5.2. Empirical Evaluation of User Experience
+NES builds upon our prior work, PEACE (Ren et al., 2025), with repo-level program understanding (Zhang et al., 2025; Tao et al., 2025; Li et al., 2024), but its reliance on live user editing data—captured exclusively via our IDE plugin—mandates an in-the-wild evaluation with internal developers.
+A direct comparison of acceptance rates is misleading due to fundamental design differences.
+Tools like CoEdPilot use a ”one-to-many” user-description-based model, which naturally produces high raw acceptance rates.
+In contrast, NES’s passive, ”one-by-one” suggestion model generates far more frequent interactions.
+To bridge this gap, we define an Equivalent Acceptance Rate (EAR) by multiplying the raw rate by the average number of lines accepted per event (1.84, based on 6 months of data).
+This normalization shows NES’s effective rates for NES-Edit and NES-Location tasks are 43.44% and 51.55%, respectively.
+These results confirm that NES offers comparable quantitative performance while providing a superior, description-free user experience.
+This experiential advantage was further validated in a comparative user study against another tool, Zeta, where NES demonstrated similar positive outcomes to those detailed in Section 4.2.
+Notably, its effectiveness becomes more pronounced as system complexity increases; in scenarios involving large classes, intricate functions, or multi-application interactions within microservices, NES more reliably completes code editing tasks than alternative methods.
+We hypothesize that this superiority stems from its core design: by capturing the fine-grained user editing trajectory, NES implicitly learns the developer’s mental model and high-level understanding of the system’s architecture—a rich signal that is absent in methods relying solely on static code snapshots.
 6. Related Work
-Transform-based Code Edits:
-Several transformer-based approaches have been proposed to extend code generation solutions to code editing tasks (Li et al., 2023; Gupta et al., 2023; Liu et al., 2024a; Wei et al., 2024; Zhang et al., 2022a).
-CODIT (Chakraborty et al., 2018) utilizes AST-based operations for code edits but faces challenges with high computation and limited semantic understanding.
-SARGAM (Liu et al., 2023) introduces a three-stage framework (Search-Generate-Modify) combining retrieval, generation, and refinement for improved accuracy and efficiency.
-CodeT5 (Wang et al., 2021) is a versatile encoder-decoder model pre-trained for various code tasks.
-Although these methods differ in how they represent edits within deep learning models, they commonly frame the edit generation problem as a translation task, which often limits their ability to deliver an improved user experience.
-General LLM-Powered Code Edits:
-In recent years, general LLMs have made significant progress in domains such as code generation (Chen et al., 2021; Zeng et al., 2025; Di et al., 2023; Jiang et al., 2024; Tian and Chen, 2023; AI, 2025), code completion (Liu et al., 2017; Alon et al., 2019; Feng et al., 2020; Liu et al., 2020; Chakraborty and Ray, 2021; Lu et al., 2022; Liang et al., 2024; Naznin and Mondal, 2023; Wang et al., 2024a; Wan et al., 2024; Wang et al., 2024b), and other code-related tasks (Liu et al., 2024b).
-Codex (Chen et al., 2021) has demonstrated the ability to generate entire functions and applications from complex natural language prompts.
-CodeFuse-13B (Di et al., 2023) is an open-sourced pre-trained Language Model (LLM) with 13 billion parameters designed for code-related tasks.
-The latest open-source model, Kimi K2 (AI, 2025), has achieved state-of-the-art (SOTA) results among open-source models on both the SWE-bench Verified (Aleithan et al., 2024) and ACeBench  (Chen et al., 2025), demonstrating its leading capabilities in code generation task.
-LLM-powered tools are well-suited for general-purpose coding tasks but often lack optimization and specialized techniques for code editing tasks, which demand stricter requirements such as low latency, high accuracy, and efficiency.
-Specialized Code Edit:
-The advent of LLMs has spurred research in code editing, leading to the development of numerous approaches tailored specifically for code editing tasks (Li et al., 2023; Gupta et al., 2023; Liu et al., 2024a; Wei et al., 2024; Jin et al., 2025).
+Code Edits with LLMs
+Early transformer-based approaches extend code generation to editing tasks (Li et al., 2023; Gupta et al., 2023; Liu et al., 2024a; Wei et al., 2024; Zhang et al., 2022b), including AST-based methods (Chakraborty et al., 2018), retrieval-augmented frameworks (Liu et al., 2023), and versatile encoder-decoder models such as CodeT5 (Wang et al., 2021).
+Recent LLMs have made significant progress in code generation (Chen et al., 2021; Zeng et al., 2025; Di et al., 2023; Codefuse et al., 2025; Jiang et al., 2024; Tian and Chen, 2023; AI, 2025), code completion (Liu et al., 2017; Alon et al., 2019; Feng et al., 2020; Liu et al., 2020; Chakraborty and Ray, 2021; Lu et al., 2022; Liang et al., 2024; Naznin and Mondal, 2023; Wang et al., 2024a; Wan et al., 2024; Wang et al., 2024b), and other code-related tasks (Liu et al., 2024b; Aleithan et al., 2024; Chen et al., 2025), with Codex (Chen et al., 2021) demonstrating the ability to generate entire functions and applications from complex natural language prompts.
+However, these tools often lack optimization for code editing’s strict requirements of low latency, high accuracy, and efficiency.
+Specialized Code Edit
+The advent of LLMs has spurred research in code editing, leading to the development of numerous approaches tailored specifically for code editing tasks (Li et al., 2023; Gupta et al., 2023; Liu et al., 2024a; Wei et al., 2024; Jin et al., 2025; Wang et al., 2025).
 InstructCoder (Li et al., 2023) introduces a dataset of 114,000 instruction-based triplets for LLM tuning, enhancing code editing accuracy for open-source models.
 However, it requires precise developer input, adding cognitive overhead.
-GRACE (Gupta et al., 2023) conditions LLMs on prior edits, significantly improving performance for edits distant from the current focus by capturing implicit developer intent.
-CoEdPilot (Liu et al., 2024a) combines historical edits, project context, and feedback to enhance code edit recommendations.
-It improves practicality but depends heavily on developer input and has high computational costs.
-Coeditor (Wei et al., 2024) uses repository-level changes and edit history to boost performance in single- and multi-turn editing.
-However, it may miss fine-grained edit patterns by focusing on repository diffs.
-Unlike these models that rely on developer prompts, NES takes a proactive approach by predicting edit locations and content.
-This enables a seamless, instruction-free workflow, fostering predictive collaboration.
-7. Conclusion and Future Work
-This paper presents NES, an LLM-powered framework for instruction-free and low-latency code editing assistance in IDEs.
-Unlike traditional approaches requiring explicit natural language instructions, NES leverages historical editing patterns to deliver efficient, context-aware support.
-It features two key components: the NES-Location Model, which predicts the next likely code modification location, and the NES-Edit Model, which generates contextually relevant code change suggestions based on editing history.
-We evaluate NES using extensive real-world datasets from both open-source and industrial projects, highlighting its practical effectiveness.
-Experimental results show that the NES-Location Model achieves 75.6% accuracy in predicting the next edit location, while NES achieves 75.6% and 81.6% accuracy in two tasks of predicting next edit locations, alongside 91.36% ES and 27.7% EMR for intent-aligned edits, outperforming SOTA models.
-Also, our open-sourced SFT and DAPO datasets have been demonstrated to enhance the performance of open-source CodeLLMs.
-Currently, NES operates within a single file, leveraging granular editing history in a localized context.
-Future work will enhance NES to handle multi-file code dependencies, enabling neural edit suggestions across interconnected files instead of single-file contexts.
+CoEdPilot (Liu et al., 2024a) enhances code edit recommendations using historical edits and project context, but relies heavily on developer input and incurs high computational costs.
+Unlike prompt-based models, NES proactively predicts edit locations and content, enabling a seamless, instruction-free workflow.
+7. Conclusion
+We present NES, an LLM framework for proactive, instruction-free code editing that learns from user editing trajectories.
+It achieves state-of-the-art performance (91.36% Edit Similarity, 27.7% Exact Match), serves 20,000+ developers at Ant Group with under 250ms latency and Acceptance Rates of 43.44% and 51.55%, and publicly releases its training datasets.